### v4

# Uniform convergence of the smooth calibration error and its relationship with functional gradient

Abstract
Calibration is a critical requirement for reliable probabilistic prediction, especially in high-risk applications. However, the theoretical understanding of which learning algorithms can simultaneously achieve high accuracy and good calibration remains limited, and many existing studies provide empirical validation or a theoretical guarantee in restrictive settings. To address this issue, in this work, we focus on the smooth calibration error (CE) and provide a uniform convergence bound, showing that the smooth CE is bounded by the sum of the smooth CE over the training dataset and a generalization gap. We further prove that the functional gradient of the loss function can effectively control the training smooth CE. Based on this framework, we analyze three representative algorithms: gradient boosting trees, kernel boosting, and two-layer neural networks. For each, we derive conditions under which both classification and calibration performances are simultaneously guaranteed. Our results offer new theoretical insights and practical guidance for designing reliable probabilistic models with provable calibration guarantees.
1 Introduction
Probabilistic prediction plays a central role in many real-world applications, such as healthcare [27], weather forecasting [42], and language modeling [43], where uncertainty estimates are often critical. Since ensuring the reliability of such predictions has become a key challenge in modern machine learning, calibration [Dawid1982, foster1998asymptotic]—which requires that predicted probabilities align with the actual frequency of the true label—has attracted increasing attention. Calibration is a relatively weak condition that can be satisfied by simple models. However, it is frequently violated in practice. Notably, guo2017calibration demonstrated that many modern deep learning models can be significantly miscalibrated, which has since drawn increasing attention to calibration within the machine learning community. Although various regularization techniques [pmlr-v80-kumar18a, karandikar2021soft, NEURIPS2022_33d6e648, marx2023calibration] and recalibration methods [zadrozny2001obtaining, guo2017calibration, gupta21b, NEURIPS2019_8ca01ea9] have been proposed to improve calibration, most of them are either empirically evaluated without a theoretical guarantee or lead to a trade-off between calibration and sharpness [Kuleshov2015], degrading predictive accuracy. As a result, it remains unclear which learning algorithms can train well-calibrated predictors without sacrificing accuracy.
One classical approach to this problem is the minimization of proper loss functions, such as the squared loss or the cross-entropy loss [schervish1989general, buja2005loss]. These losses are minimized in expectation by the true conditional probability and thus have a natural connection to calibration. However, in practice, constraints on model classes and suboptimal optimization may prevent proper losses from yielding calibrated predictions. To address this issue, in a recent theoretical work, the notion of a post-processing gap [blasiok2023when] has been introduced, which quantifies the potential improvement of the loss function achievable by applying Lipschitz-continuous transformations to the model outputs. This framework provides an optimization-theoretic perspective on calibration. Nevertheless, analyses based on this concept typically assume access to infinite data (i.e., population-level risk) [blasiok2023loss, pmlr-v202-globus-harris23a], which limits their applicability to real-world algorithms based on finite training samples. Furthermore, only a few theoretical results explicitly connect calibration error with concrete learning algorithms [hansen2024multicalibration, blasiok2023loss]. As a result, a key question remains unresolved: What types of learning algorithms can achieve both high predictive accuracy and strong calibration guarantees in practice?
To address this gap, we focus on the smooth calibration error (CE) [blasiok2023unify, foster2018smooth, kakade2008deterministic], a recently proposed calibration metric for binary classification that has favorable theoretical properties. Using smooth CE, we contribute to answering the above question through two main theoretical advances. First, we derive a uniform convergence bound for smooth CE, showing that the population-level smooth CE can be bounded by the smooth CE over the training dataset plus a generalization gap (Section 3). Second, we prove that such training smooth CE can be bounded by the norm of the functional gradient (or its approximation) of the loss evaluated on training data, providing a principled criterion for optimizing calibration (Section 4). Taken together, these results show that algorithms that achieve small functional gradients while regularizing model complexity can obtain small smooth CE.
We apply our theoretical framework to three representative algorithms for the binary classification closely tied to functional gradients: gradient boosting trees (GBTs), kernel boosting, and two-layer neural networks (NNs). For GBTs, we provide the theoretical guarantee for their calibration performance, showing that the smooth CE decreases with iteration (Section 4.1). For kernel boosting, we analyze the trade-off between optimization and model complexity induced by the reproducing kernel Hilbert space (Section 4.2). For two-layer NNs, we leverage their connection to the neural tangent kernel to study calibration behavior on the basis of function space optimization (Section 4.3). Furthermore, for all these algorithms, by introducing a margin-based assumption, we derive sufficient conditions on the sample size and the number of iterations required to simultaneously achieve $\epsilon$-level smooth CE and misclassification rate under appropriately chosen hyperparameters.
In summary, in this work, we develop a unified theoretical framework for the smooth CE that integrates a uniform convergence analysis with a functional-gradient-based characterization of training dynamics. Our results establish new theoretical foundations for understanding and improving calibration in modern learning algorithms.
2 Preliminaries
In this section, we introduce proper losses and the smooth CE.
2.1 Problem setting of binary classification
We denote random variables by capital letters (e.g., $X$) and deterministic values by lowercase letters (e.g., $x$). The Euclidean inner product and norm are denoted by $\cdot$ and $\|\cdot\|$, respectively.
We consider a binary classification problem in the supervised setting. Let $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ denote the data domain, where $\mathcal{X}\subset\real^{d}$ is the input space and $\mathcal{Y}=\{0,1\}$ is the label space. We assume that data points are sampled from an unknown data-generating distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$. Let $\mathcal{F}$ be a class of predictors $f:\mathcal{X}\to[0,1]$ approximating the ground-truth conditional probability $f^{*}(x)=\mathbb{E}[Y|x]$.
To evaluate the performance of the predictor, we use a loss function. Since the task is binary classification, the label $Y$ follows a Bernoulli distribution, i.e., $Y\sim\text{Ber}(v)$ for some $v\in[0,1]$, where ${\mathrm{Pr}}(Y=1)=v$. A loss function is defined as $\ell:[0,1]\times\mathcal{Y}\to\mathbb{R}$, with evaluation given by $\ell(v,y)$. Our goal is not only to achieve high classification accuracy but also to assess the quality of the predicted probabilities. In this context, proper loss functions play a central role [Gneiting2007]. The loss $\ell$ is called proper if, for any $v,v^{\prime}\in[0,1]$, the following holds: $\mathbb{E}_{Y\sim\text{Ber}(v)}[\ell(v,Y)]\leq\mathbb{E}_{Y\sim\text{Ber}(v)}[\ell(v^{\prime},Y)]$.
For clarity, we focus on two representative proper losses: the squared loss $\ell_{\mathrm{sq}}(v,y)\coloneqq(y-v)^{2}$ and the cross-entropy loss $\ell_{\mathrm{ent}}(v,y)\coloneqq-y\log v-(1-y)\log(1-v)$. Given $x$ and $f$, we substitute $v=f(x)$ in these expressions. For other proper losses, see Appendix A and blasiok2023when.
2.2 Calibration metrics and loss functions
Given a distribution $\mathcal{D}$ and a predictor $f$, we say the model is perfectly calibrated if $\mathbb{E}[Y\mid f(X)]=f(X)$ holds almost surely [blasiok2023unify]. Various metrics have been proposed to quantify the deviation from perfect calibration. The most widely used metric is the expected calibration error (ECE): $\mathrm{ECE}(f):=\mathbb{E}\left[\left|\mathbb{E}[Y\mid f(X)]-f(X)\right|\right].$ Despite its popularity, ECE is difficult to estimate efficiently [arrieta2022metrics, lee2023t] and is known to be discontinuous [foster2018smooth, kakade2008deterministic]. As a more tractable alternative, we focus on the smooth CE:
Definition 1 (Smooth CE [blasiok2023unify]).
Let $\mathrm{Lip}_{L}([0,1],[-1,1])$ be the set of $L$-Lipschitz functions from $[0,1]$ to [1,1][-1,1]. Given $f$ and $\mathcal{D}$, the smooth CE is defined as
$\mathrm{smCE}(f,\mathcal{D})\coloneqq\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\mathbb{E}\left[h(f(X))\cdot(Y-f(X))\right].$ (1)
We remark that the concept of the smooth CE is also introduced in foster2018smooth, kakade2008deterministic. We also express $\mathrm{smCE}(f,\mathcal{D})=\sup_{h\in\mathrm{Lip}_{1}([0,1],[-1,1])}\left\langle h(f(X)),Y-f(X)\right\rangle_{L_{2}(\mathcal{D})}$. The smooth CE offers favorable properties such as continuity and computational tractability [blasiok2023unify, hu2024testing]. Moreover, the smooth CE provides both upper and lower bounds on the binning ECE, making it a principled surrogate for analyzing calibration. See Section 5 for details.
blasiok2023when established a connection between the smooth CE and the optimization based on the concept of the possible improvement of the function under the squared loss. Given $f$ and $\mathcal{D}$, we define the post-processing gap as
$\mathrm{pGap}(f,\mathcal{D})\coloneqq\mathbb{E}[\ell_{\mathrm{sq}}(f(X),Y)]-\inf_{h\in\mathrm{Lip}_{1}([0,1],[-1,1])}\mathbb{E}[\ell_{\mathrm{sq}}(f(X)+h(f(X)),Y)].$ (2)
Then, the following inequality holds (Theorem 2.4 in blasiok2023when):
$\mathrm{smCE}(f,\mathcal{D})^{2}\leq\mathrm{pGap}(f,\mathcal{D})\leq 2\mathrm{smCE}(f,\mathcal{D}).$ (3)
blasiok2023when also provided analogous results for the general proper scoring rule, which indicates that the potential improvement in the population risk is closely tied to calibration quality.
We now consider the setting of the cross-entropy loss. In practice, the predicted probability $f$ is often obtained by applying the sigmoid function $\sigma$ to a logit function $g:\mathcal{X}\to\mathbb{R}$, i.e., $f(x)=\sigma(g(x))=1/(1+e^{-g(x)})$. Therefore, it is more natural to consider post-processing over the logit $g$ rather than the predicted probability $f$, as many models explicitly learn logits and apply the sigmoid function only at the final prediction stage. We define a corresponding post-processing gap over the logit function with the cross-entropy loss as
$\mathrm{pGap}^{\sigma}(g,\mathcal{D})\coloneqq\mathbb{E}[\ell_{\mathrm{ent}}(\sigma(g(X)),Y)]-\inf_{h\in\mathrm{Lip}_{1}(\real,[-4,4])}\mathbb{E}[\ell_{\mathrm{ent}}(\sigma(g(X)+h(g(X))),Y)],$ (4)
where $\mathrm{Lip}_{1}(\real,[-4,4])$ defines a class of $1$-Lipschitz functions from to [4,4][-4,4] and the choice of $4$ reflects the fact that the sigmoid function is $1/4$-Lipschitz. Then, this post-processing leads to defining a smooth CE directly over the logit function:
Definition 2 (Dual smooth CE [blasiok2023unify]).
Given a logit function $g:\mathcal{X}\to\mathbb{R}$ and $f(x)=\sigma(g(x))$, the dual smooth CE is defined as
$\mathrm{smCE}^{\sigma}(g,\mathcal{D})\coloneqq\sup_{h\in\mathrm{Lip}_{1/4}(\real,[-1,1])}\left\langle h(g(X)),Y-f(X)\right\rangle_{L_{2}(\mathcal{D})}.$ (5)
Similarly to Eq. equation 3, the following relation holds:
$2\mathrm{smCE}^{\sigma}(g,\mathcal{D})^{2}\leq\mathrm{pGap}^{\sigma}(g,\mathcal{D})\leq 4\mathrm{smCE}^{\sigma}(g,\mathcal{D}).$ (6)
Finally, we note that the dual smooth CE always upper bounds the smooth CE: $\mathrm{smCE}(f,\mathcal{D})\leq\mathrm{smCE}^{\sigma}(g,\mathcal{D})$. Hence, minimizing the dual smooth CE guarantees a small value of the smooth CE.
3 Uniform convergence of the smooth CE
As shown in Eqs. equation 3 and equation 6, (dual) smooth CE is related to improvements in the population risk, and blasiok2023when utilized these relationships to analyze when algorithms can achieve a small smooth CE. However, their results are defined over population-level quantities, where the expectation by $\mathcal{D}$ is taken, making them inapplicable to practical algorithms trained on finite data points.
Since our goal is to analyze the behavior of the smooth CE under standard learning algorithms trained on finite data, it is natural to consider its empirical counterpart. Given a dataset $S_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ consisting of independently and identically distributed (i.i.d.) data points sampled from the data distribution $\mathcal{D}$, we define the empirical smooth CE [blasiok2023unify] given $S_{n}$ as
$\mathrm{smCE}(f,S_{n})\coloneqq\sup_{h\in\mathrm{Lip}_{1}([0,1],[-1,1])}\frac{1}{n}\sum_{i=1}^{n}h(f(X_{i}))\cdot(Y_{i}-f(X_{i})).$ (7)
We also express this as $\mathrm{smCE}(f,S_{n})=\sup_{h\in\mathrm{Lip}_{1}([0,1],[-1,1])}\left\langle h(f(X)),Y-f(X)\right\rangle_{L_{2}(S_{n})}$. We similarly define $\mathrm{smCE}^{\sigma}(g,S_{n})$ as empirical counterparts of $\mathrm{smCE}^{\sigma}(g,\mathcal{D})$.
We consider two datasets: a training set $S_{\mathrm{tr}}=\{Z_{i}\}_{i=1}^{n}\sim\mathcal{D}^{n}$, used to learn the predictor $f$, and an independent test set $S_{\mathrm{te}}=\{Z^{\prime}_{i}\}_{i=1}^{n}\sim\mathcal{D}^{n}$. We call $\mathrm{smCE}(f,S_{\mathrm{tr}})$ as the training smooth CE, $\mathrm{smCE}(f,S_{\mathrm{te}})$ as the test smooth CE, and $\mathrm{smCE}(f,\mathcal{D})$ as the population smooth CE. Given $S_{\mathrm{tr}}$, we are interested in evaluating $|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$. blasiok2023unify has shown that
$|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{te}})|=\mathcal{O}_{p}(1/\sqrt{n}).$ (8)
Therefore, we need to evaluate $|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$, which we refer to as the smooth CE generalization gap. Combining this gap with Eq. equation 8, we obtain the desired bound. To evaluate the smooth CE generalization gap, we use the covering number bound. Suppose that $\mathcal{F}$ is equipped with the metric $\|\cdot\|_{\infty}$. Let $\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$ be an $\epsilon$-cover with metric $\|\cdot\|_{\infty}$ of $\mathcal{F}$, with the cardinality $N(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$. Then, for any $f\in\mathcal{F}$, there exists $\tilde{f}\in\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{\infty})$ such that $\|f-\tilde{f}\|_{\infty}\leq\epsilon$.
We now present our first main result: (All proofs for this section are provided in Appendix C.)
Theorem 1.
Let $S_{\mathrm{te}}\sim\mathcal{D}^{n}$ and $S_{\mathrm{tr}}\sim\mathcal{D}^{n}$ be independent test and training datasets. Then, for any $\delta>0$, with probability at least $1-\delta$ over the draw of $S_{\mathrm{te}}$ and $S_{\mathrm{tr}}$, we have:
$\!\sup_{f\in\mathcal{F}}|\mathrm{smCE}(f,S_{\mathrm{te}})\!-\!\mathrm{smCE}(f,S_{\mathrm{tr}})|\!\leq\!\inf_{\epsilon\geq 0}8\epsilon\!+\!24\!\int_{\epsilon^{\prime}}^{1}\!\sqrt{\frac{\ln N(\epsilon^{\prime},\mathcal{F},\|\cdot\|_{\infty})}{n}}\!d\epsilon^{\prime}\!+\!2\sqrt{\frac{\log\delta^{-1}}{n}}.$ (9)
Since the smooth CE involves evaluating the composite function $h(f(X))$, standard chaining bounds would introduce complexity over the composite class $\mathrm{Lip}_{1}([0,1],[-1,1])\circ\mathcal{F}$. However, by leveraging the smoothness property of smooth CE, the above bound does not include the complexity of Lipschitz functions. Combining Theorem 1 with Eq. equation 8, we obtain the bound for the population smooth CE.
Corollary 1.
Under the same assumptions as in Theorem 1, there exist a universal constant $C_{1}$ such that with probability at least $1-\delta$ over the draw of $S_{\mathrm{tr}}$, the following holds for all $f\!\in\!\mathcal{F}$:
$|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ $\leq\frac{C_{1}}{\sqrt{n}}+\inf_{\epsilon\geq 0}8\epsilon\!+\!24\!\int_{\epsilon^{\prime}}^{1}\!\sqrt{\frac{\ln N(\epsilon^{\prime},\mathcal{F},\|\cdot\|_{\infty})}{n}}d\epsilon^{\prime}\!+\!3\sqrt{\frac{\log\frac{3}{\delta}}{n}}.$ (10)
The first term of the right-hand side represents the complexity of the Lipschitz function class.
In certain cases, the Rademacher complexity offers a more interpretable characterization. The empirical and expected Rademacher complexities of a function class $\mathcal{F}$ are defined as $\hat{\mathfrak{R}}_{S}(\mathcal{F})\coloneqq\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}f(x_{i})$ , and $\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})\coloneqq\mathbb{E}_{\mathcal{D}^{n}}[\hat{\mathfrak{R}}_{S}(\mathcal{F})]$, respectively.
Theorem 2.
Under the same assumptions as in Theorem 1, there exist a universal constant $C_{2}$ such that with probability at least $1-\delta$ over the draw of $S_{\mathrm{tr}}$, the following holds:
$\sup_{f\in\mathcal{F}}|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ $\leq\frac{C_{2}}{\sqrt{n}}+4\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (11)
The first term on the right-hand side reflects the complexity of $\mathrm{Lip}_{1}([0,1],[-1,1])$, while the second corresponds to $\mathcal{F}$. A direct application of uniform convergence theory would yield a bound in terms of $\mathfrak{R}_{\mathcal{D},n}(\mathrm{Lip}_{1}([0,1],[-1,1])\circ\mathcal{F})$, which cannot be reduced to $\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})$ using the standard contraction lemma. Our analysis circumvents this difficulty by employing a covering argument for Lipschitz functions. Similarly, Rademacher complexity can be used to bound $|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$. However, such a bound depends on the complexity of the Lipschitz function, making it suboptimal relative to Theorem 1. See Appendix D for a detailed comparison between Corollary 1 and Theorem 2.
In conclusion, by jointly controlling the complexity of the hypothesis class and reducing the training smooth CE, we guarantee a small population smooth CE. In the next section, we show how to achieve a small training smooth CE via functional gradients.
4 Controlling the smooth CE via functional gradient
In this section, we discuss how we can minimize the training smooth CE. From Eqs. equation 3 and equation 6, (dual) smooth CE can be characterized via the (dual) post-processing gap, which describes the potential improvement over functions. Building on this perspective, we show that both smooth and dual smooth CEs can be further characterized using functional gradients. Since we focus on the functional gradients over the training dataset, they are just finite-dimensional vectors. We provide the precise connection between functional gradients and the population smooth CE in Appendix H.
For the squared loss, the gradient for the predicted probability is $\nabla_{f}\ell_{\mathrm{sq}}(f,y)=f-y$, and for the cross-entropy loss, the gradient for the logit is $\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g),y)=\sigma(g)-y$. Using these, we can write the smooth and dual smooth CEs on the training dataset $S_{\mathrm{tr}}$ as
$\mathrm{smCE}(f,S_{\mathrm{tr}})$ $=\sup_{h\in\mathrm{Lip}_{1}([0,1],[-1,1])}\left\langle h(f(X)),-\nabla_{f}\ell_{\mathrm{sq}}(f(X),Y)\right\rangle_{L_{2}(S_{n})},$ (12)
$\mathrm{smCE}^{\sigma}(g,S_{\mathrm{tr}})$ $=\sup_{h\in\mathrm{Lip}_{1/4}(\real,[-1,1])}\left\langle h(g(X)),-\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X)),Y)\right\rangle_{L_{2}(S_{n})}.$ (13)
The connection between functional gradients and the smooth CE can be extended to general proper scoring rules using the post-processing gaps, see Appendix H for the details.
Therefore, to effectively control the training smooth CE, it is natural to consider algorithms that focus on functional gradients. From this perspective, in the following subsections, we investigate how the smooth CE interacts with the training dynamics of three widely used models: gradient boosting tree, kernel boosting in a reproducing kernel Hilbert space (RKHS), and two-layer neural networks. These models can all be interpreted on the basis of functional gradients, and our theoretical framework enables a unified understanding of their calibration behavior under different forms of regularization.
In this section, we express the empirical risk as $L_{n}(g)=\frac{1}{n}\sum_{i=1}^{n}\ell_{\mathrm{ent}}(\sigma(g(X_{i})),Y_{i})$ and the functional gradient over the training dataset as $\nabla_{g}L_{n}(g)=(\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{1})),Y_{1})),\dots,\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{n})),Y_{n}))\in\real^{n}$.
4.1 Case study I: Gradient Boosting Tree
Gradient boosting [Friedmangreedy2001] is a widely used method for constructing predictive models by iteratively adding base learners to minimize a loss function. It generates a sequence of functions $\{g^{(t)}\}_{t=0}^{T}$ via the updates, for $t=0,\dots,T-1$:
$g^{(t+1)}(x)=g^{(t)}(x)-w_{t}\psi_{t}(x),$ (14)
where $w_{t}>0$ is the stepsize and $\psi_{t}$ approximates the functional gradient of the empirical loss [mason1999boosting]. Since using the exact functional gradient often leads to overfitting, $\psi_{t}$ is restricted to a predefined function class $\Psi$, which provides implicit regularization. We impose the following mild assumption on $\Psi$, which is readily satisfied by common choices such as regression trees.
Assumption 1.
The set of real-valued functions $\Psi$ satisfies: for every $\psi\in\Psi$, the negation $\psi$ also belongs to $\Psi$; $\sup_{\psi\in\Psi}\|\psi\|_{\infty}\leq B$ for some $B\geq 1$; and the constant function 0 is included in $\Psi$.
To obtain the update direction, we iteratively solve $\psi^{(t)}=\operatorname*{arg\,min}_{\psi\in\Psi}L_{n}\big(g^{(t)}-w_{t}\psi(x)\big)$. Many boosting methods approximate this by a quadratic upper bound, leading to the squared loss
$\psi_{t}=\operatorname*{arg\,min}_{\psi\in\Psi}\|Mw_{t}\psi-\nabla_{g}L_{n}(g^{(t)})\|^{2}_{L_{2}(S_{n})}=\frac{1}{n}\sum_{i=1}^{n}|Mw_{t}\psi(X_{i})-\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{i})),Y_{i}))|^{2}.$ (15)
where $M$ is the smoothness parameter of the loss (e.g., $M=1/4$ for cross-entropy).
Although gradient boosting is often observed to be well calibrated in practice [niculescu2005obtaining, wenger2020non], its theoretical guarantees remain less understood. Our framework provides a natural way to analyze this. Following prior analysis on boosting, we impose a standard margin assumption:
Assumption 2 (telgarsky2013margins).
Let $\Delta_{n}=\{q\in\real_{+}^{n}\mid\sum_{i=1}^{n}q_{i}=1\}$ denote the $n$-dimensional probability simplex. Given a dataset $S_{\mathrm{tr}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, there exists $\gamma>0$ such that for every $q\in\Delta_{n}$, there exists $\psi\in\Psi$ satisfying $\sum_{i=1}^{n}q_{i}(2Y_{i}-1)\psi(X_{i})\geq\gamma B$.
Here, $2Y_{i}-1$ maps the binary label $Y_{i}\in\{0,1\}$ to $\pm 1$. This condition guarantees that for any weighted sample distribution $q\in\Delta_{n}$, one can choose a base learner with nontrivial correlation to the labels. Such margin assumptions are also standard in classification theory [wei2018margin]. Under this setup, we present our main result for the training dual smooth CE:
Theorem 3.
Given $S_{\mathrm{tr}}$, under Assumptions 1 and 2, if $T\geq 2$ with constant stepsize $w_{t}=w$, the averaged predictor $\bar{g}^{(T)}=\frac{1}{T}\sum_{t=0}^{T-1}g^{(t)}$ satisfies
$\mathrm{smCE}^{\sigma}\big(\bar{g}^{(T)},S_{n}\big)\leq\frac{L_{n}(g^{(0)})}{\gamma BwT}+\frac{wB}{8\gamma}.$ (16)
Proof outline.
By definition, the smooth CE is bounded by the $L_{1}$ norm of the functional gradient:
$\mathrm{smCE}^{\sigma}(g,S_{\mathrm{tr}})\leq\frac{1}{n}\sum_{i=1}^{n}|\nabla_{g}\ell_{\mathrm{ent}}(g(X_{i}),Y_{i})|=\|\nabla_{g}\ell_{\mathrm{ent}}(g(X),Y)\|_{L_{1}(S_{n})}.$ (17)
Under Assumption 2, this norm can be further bounded as
$\|\nabla_{g}\ell_{\mathrm{ent}}(g(X),Y)\|_{L_{1}(S_{n})}\leq\frac{1}{\gamma B}\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}+\frac{w}{8\gamma}B$ (18)
as shown in Lemma 4 in Appendix E. The inner product $\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}$ is then bounded based on the standard convex optimization techniques. See Appendix E.3 for the complete proof. ∎
Thus, choosing $w=\mathcal{O}(1/\sqrt{T})$ ensures that the training smooth CE converges to 0 at rate $\mathcal{O}(1/\sqrt{T})$.
Next, we establish a guarantee for the population smooth CE by applying Theorem 2. This requires specifying $\Psi$ to evaluate the complexity of $\mathcal{F}$. We consider gradient boosting trees [friedman2002stochastic, hastie2005elements], where $\Psi$ is the class of binary regression trees. A binary regression tree of depth $m$ partitions the input space $\mathbb{R}^{d}$ into at most $J\leq 2^{m}$ disjoint regions: $\mathbb{R}^{d}=\bigcup_{j=1}^{J}R_{j},\text{with }R_{j}\cap R_{k}=\emptyset\text{ for }j\neq k.$ Each region $R_{j}$ is assigned a constant $c_{j}\in\mathbb{R}$, yielding a piecewise constant function: $\psi_{\theta}(x)=\sum_{j=1}^{J}c_{j}\cdot\mathbbm{1}_{\{x\in R_{j}\}}$ with parameters $\theta=\{c_{j},R_{j}\}_{j=1}^{J}$. The complete GBT algorithm is provided in Appendix E. Under these settings, we obtain the following result:
Corollary 2.
Under the same assumptions as Theorem 3, there exist universal constants $\{C_{i}\}$s such that, with probability at least $1-\delta$ over the draw of $S_{\mathrm{tr}}$, we have
$\mathrm{smCE}(\sigma(\bar{g}^{(T)}),\mathcal{D})\leq\frac{L_{n}(g^{(0)})}{\gamma BwT}+\frac{wB}{8\gamma}+\frac{C_{2}}{\sqrt{n}}+C_{3}wT\sqrt{\frac{2^{m}\log nd}{n}}+2\sqrt{\frac{\log(2/\delta)}{n}}.$ (19)
We find that increasing the number of steps $T$ reduces the training smooth CE, but also enlarges the function class, as the Rademacher complexity grows at rate $\mathcal{O}(wT)$. This highlights a trade-off between lowering training smooth CE and controlling model complexity. Corollary 3 shows that with appropriate hyperparameters, one can attain any target precision $\epsilon$ for the smooth CE.
In addition, bounding the norm of the functional gradient provides a generalization guarantee for test accuracy [nitanda18a]. For completeness, we include in Appendix E a formal upper bound on the misclassification rate, $P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]$.
Combining this with Theorem 2, we conclude that with suitable choices of $T$ and $n$, both the smooth CE and the misclassification rate can be controlled to within any target precision $\epsilon$.
Corollary 3.
Under assumptions in Corollary 2, for any $\epsilon>0$, if the hyperparameters satisfy:
$T=\Omega(\gamma^{-2}\epsilon^{-2}),\quad w=\Theta(\gamma^{-1}\epsilon^{-1}T^{-1}),\quad n=\tilde{\Omega}(\gamma^{-2}\epsilon^{-4}),$ (20)
then, with probability at least $1-\delta$, the averaged predictor $\bar{g}^{(T)}$ satisfies $\mathrm{smCE}(\sigma(\bar{g}^{(T)}),\mathcal{D})\leq\epsilon$ and $P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq\epsilon$.
[…]
4.2 Case study II: Kernel Boosting
We next consider kernel boosting [wei2017early], where the functional gradient is approximated by functions in an RKHS. Let $(\mathcal{H},\langle\cdot,\cdot\rangle_{\mathcal{H}})$ be an RKHS associated with the kernel $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$. The functional gradient over the training dataset is approximated as
$\max_{\phi\in\mathcal{H},\|\phi\|_{\mathcal{H}}\leq 1}\left\langle\nabla_{g}L_{n}(g^{(t)}),\phi\right\rangle_{L_{2}(S_{n})}=\frac{\mathcal{T}_{k}\nabla_{g}L_{n}(g)}{\|\mathcal{T}_{k}\nabla_{g}L_{n}(g)\|_{\mathcal{H}}},$ (21)
where $\mathcal{T}_{k}\nabla_{g}L_{n}(g)\coloneqq\frac{1}{n}\sum_{i=1}^{n}k(X_{i},\cdot)\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g(X_{i})),Y_{i}))$ is the empirical kernel operator. In kernel boosting, we set $h_{t}=-\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})$ in Eq. equation 14 and use the update rule:
$g^{(t+1)}=g^{(t)}-w_{t}\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)}).$ (22)
To analyze this, we introduce the following normalized margin assumption:
Assumption 3.
Given a dataset $S_{\mathrm{tr}}$, there exists a function $\phi\in\mathcal{H}$ and a constant $\gamma>0$ such that for all $(X_{i},Y_{i})\in S_{\mathrm{tr}}$, $\left(2Y_{i}-1\right)\phi(X_{i})\geq\gamma$.
This assumption is essentially equivalent to Assumption 2 in the GBT analysis. Setting $q$ in Assumption 2 as a unit vector yields Assumption 3. Conversely, taking convex combinations of $\left(2Y_{i}-1\right)\phi(X_{i})\geq\gamma$ in Assumption 3 recovers Assumption 2.
Under Assumption 3, the $L_{1}$ norm of the functional gradient is bounded as follows:
$\|\nabla_{g}\ell_{\mathrm{ent}}(g(X),Y)\|_{L_{1}(S_{n})}=\frac{1}{n}\sum_{i=1}^{n}|\nabla_{g}\ell_{\mathrm{ent}}(g(X_{i}),Y_{i})|\leq\frac{1}{\gamma}\left\|\mathcal{T}_{k}\nabla_{g}L_{n}(g)\right\|_{\mathcal{H}}.$ (23)
All proofs for this section are provided in Appendix F. Combining this with Eq. equation 17, we can control the smooth CE via $\left\|\mathcal{T}_{k}\nabla_{g}L_{n}(g)\right\|_{\mathcal{H}}$. Using this relation, we obtain the following result.
Theorem 4.
Suppose Assumption 3, $\sup_{x,x^{\prime}\in\mathcal{X}}k(x,x^{\prime})\leq\Lambda$, and $\|g^{(0)}\|_{\mathcal{H}}\leq\Lambda^{\prime}$ hold. When using the constant stepsize $w_{t}=w$ which satisfies $w<4/\Lambda$, the average of function $\bar{g}^{(T)}=\frac{1}{T}\sum_{t=0}^{T-1}g^{(t)}$ satisfies
$\mathrm{smCE}^{\sigma}(\bar{g}^{(T)},S_{\mathrm{tr}})\leq\frac{1}{\gamma}\sqrt{\frac{L_{n}(g^{(0)})}{wT}}.$ (24)
Additionally, there exist universal constants $\{C_{i}\}$s such that with probability at least $1-\delta$ over the draw of $S_{\mathrm{tr}}$, we have:
$\mathrm{smCE}(\sigma(\bar{g}^{(T)}),\mathcal{D})\leq\frac{1}{\gamma}\sqrt{\frac{L_{n}(g^{(0)})}{wT}}\!+\frac{C_{2}}{\sqrt{n}}\!+C_{4}(\Lambda^{\prime}+\sqrt{2wTL_{n}(g^{(0)})})\sqrt{\frac{\Lambda}{n}}\!+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (25)
Although covering number bounds require assumptions on kernel eigenvalue decay, here we use the Rademacher complexity bound from Theorem 2, which yields a simpler result.
We observe that increasing the number of steps $T$ decreases the training smooth CE but simultaneously enlarges the function class, since the Rademacher complexity of predictors (the third term on the right-hand side) grows with the norm of $\bar{g}^{(t)}$ at rate $\mathcal{O}(\sqrt{wT})$. This highlights a trade-off between reducing training smooth CE and controlling model complexity.
Since the norm of the functional gradient also upper bounds the misclassification rate, we show—analogous to Corollary 3—that suitable hyperparameter choices guarantee any target precision $\epsilon$ for both the smooth CE and the misclassification rate.
Corollary 4.
Suppose assumptions in Theorem 4 hold. For any $\epsilon>0$, if the hyperparameters satisfy:
$T=\Omega(\gamma^{-2}\epsilon^{-2}),\quad w=\Theta(\gamma^{-2}\epsilon^{-2}T^{-1}),\quad n=\tilde{\Omega}(\gamma^{-2}\epsilon^{-4}),$ (26)
then, with probability at least $1-\delta$, the average of function $\bar{g}^{(T)}$ satisfies $\mathrm{smCE}(\sigma(\bar{g}^{(T)}),\mathcal{D})\leq\epsilon$ and $P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq\epsilon$.
Corollary 4 establishes that both $\epsilon$-smooth CE and $\epsilon$-classification error can be achieved using $\mathcal{O}(1/\epsilon^{2})$ iterations of kernel boosting. This result is grounded in the observation that bounding the functional gradient norm leads to a small smooth CE and a misclassification rate.
4.3 Case study III: Two-layer neural network
The next example is the two-layer neural network (NN). While NNs are typically trained by gradient descent on their parameters, recent work [nitanda2019gradient] shows that under certain hyperparameter settings—such as output scaling and a sufficiently large number of neurons—two-layer NNs behave similarly to kernel boosting with the neural tangent kernel (NTK).
We adopt the setting of nitanda2019gradient. Define the logit function $g_{\theta}:\mathcal{X}\rightarrow\mathbb{R}$ by a two-layer NN $g_{\theta}(x)=\frac{1}{m^{\beta}}\sum_{r=1}^{m}a_{r}\phi(\theta_{r}\cdot x)$, where $m>0$ is the number of hidden units, $\beta\in[0,1]$ is a scaling exponent, and $\phi:\mathbb{R}\to\mathbb{R}$ is a smooth activation function (e.g., sigmoid, tanh). The weights $\{a_{r}\}_{r=1}^{m}\in\{-1,1\}^{m}$ are fixed, while the parameters $\theta=\{\theta_{r}\}_{r=1}^{m}$ with $\theta_{r}\in\mathbb{R}^{d}$ are updated via gradient descent: $\theta^{(t+1)}=\theta^{(t)}-w\nabla_{\theta}L_{n}(g_{\theta^{(t)}})$ with constant stepsize $w>0$.
We now give an informal statement of Theorem 2 from nitanda2019gradient, which provides guarantees for the functional gradient. Assume: (1) the activation function is smooth; (2) the initial parameters are sampled from a sub-Gaussian distribution; (3) the NTK-transformed data are separable with margin $\gamma$; (4) the stepsize is sufficiently small; and (5) the number of hidden units $m$ is large enough. Then, for $T\leq\frac{m\gamma^{2}K_{3}}{w}$, there exist constants $K_{1}$, $K_{2}$, and $K_{3}$ such that, for all $\beta\in[0,1]$, with probability at least $1-\delta$ over the random initialization, the following bound holds:
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla_{g}\ell_{\mathrm{ent}}(g_{\theta^{(t)}}(X),Y)\|^{2}_{L_{1}(S_{n})}\leq\frac{K_{1}}{\gamma^{2}T}\left(\frac{m^{2\beta-1}}{w}+K_{2}\right).$ (27)
See Appendix G for the formal statement. Then, similar to Theorem 4, we have
$\min_{t\in\{0,\dots,T-1\}}\mathrm{smCE}^{\sigma}(g_{\theta^{(t)}},S_{\mathrm{tr}})\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathrm{smCE}^{\sigma}(g_{\theta^{(t)}},S_{\mathrm{tr}})\leq\sqrt{\frac{K_{1}}{\gamma^{2}T}\left(\frac{m^{2\beta-1}}{w}+K_{2}\right)}.$ (28)
Combining the Rademacher complexity estimate of nitanda2019gradient with Theorem 2, we derive upper bounds on the population smooth CE and misclassification rate, analogous to the kernel boosting setting (see Appendix G for details). As in Theorem 4, the bound reveals a trade-off in $T$: increasing $T$ reduces the training smooth CE but enlarges the complexity term, which grows at rate $\mathcal{O}(\sqrt{wT})$. Similar to Corollary 4, we further show that with suitable hyperparameters, both $\epsilon$-smooth CE and $\epsilon$-classification error can be guaranteed.
Corollary 5 (Informal).
Suppose the same assumptions as those for Eq. equation 27 hold. If for any $\epsilon>0$, the hyperparameters satisfy one of the following:
(i) $\beta\in[0,1)$, $m=\Omega(\gamma^{\frac{-2}{1-\beta}}\epsilon^{\frac{-1}{1-\beta}})$, $T=\Omega(\gamma^{-2}\epsilon^{-2})$, $w=\Theta(\gamma^{-2}\epsilon^{-2}T^{-1}m^{2\beta-1})$, $n=\tilde{\Omega}(\gamma^{-2}\epsilon^{-4})$,
(ii) $\beta=0$, $m=\Theta\left(\gamma^{-2}\epsilon^{-3/2}\log(1/\epsilon)\right)$, $T=\Theta\left(\gamma^{-2}\epsilon^{-1}\log^{2}(1/\epsilon)\right)$, $w=\Theta(m^{-1})$, $n=\tilde{\Omega}(\epsilon^{-2})$, then with probability at least $1-\delta$, gradient descent with the stepsize $w$ finds a parameter $\theta^{(t)}$ satisfying $\mathrm{smCE}(\sigma(g_{\theta^{(t)}}),S_{\mathrm{tr}})\leq\epsilon$ and $\mathbb{P}_{(X,Y)\sim\mathcal{D}}\left[(2Y-1)g_{\theta^{(t)}}(X)\leq 0\right]\leq\epsilon$ within $T$ iterations.
See Appendix G for the formal statement with explicit constants. This result shows that both $\epsilon$-smooth CE and $\epsilon$-classification error can be achieved within $\mathcal{O}(1/\epsilon^{2})$ or $\mathcal{O}(1/\epsilon)$ iterations, assuming the NTK-transformed data are separable with margin $\gamma$. Between the two hyperparameter settings, (ii) attains the same error target with fewer iterations $T$ by increasing the number of hidden units $m$, and further yields improved complexity compared with Corollary 4 for kernel boosting. Appendix I provides empirical results illustrating how smooth CE and accuracy vary with $T$ and $n$.
5 Related work
Various metrics have been proposed to quantify deviations from perfect calibration, including ECE [guo2017calibration, rahaman2021uncertainty, minderer2021revisiting]. The binning ECE has been widely used to estimate the conditional expectation $\mathbb{E}[Y|f(X)]$ in the ECE, which partitions the probability interval into discrete bins. Despite its popularity, binning ECE is sensitive to hyperparameters such as the number of bins, and it lacks both consistency and smoothness as a distance metric [kumar2019verified, nixon2019measuring, minderer2021revisiting]. A broader comparison of ECE variants is given by gruber2022better. blasiok2023unify systematically studied calibration distances and formalized the true distance from calibration as a rigorous notion of deviation from perfect calibration. As an efficiently estimable surrogate, they proposed the smooth CE, building on foster2018smooth, kakade2008deterministic. More recently, hu2024testing showed that smooth CE can be efficiently evaluated in practice. Importantly, blasiok2023unify proved that smooth CE provides both upper and lower bounds on binning ECE, making it a theoretically sound proxy. Hence, the results established here for smooth CE also yield implications for binning ECE; see Appendix B for details.
To achieve a well-calibrated prediction with a theoretical guarantee, many prior works focus on recalibration, such as binning-based recalibration methods [gupta2020, gupta21b, kumar2019verified, sun2023minimumrisk, NEURIPS2024_futami]. However, it has been reported that such post-processing loses the sharpness of the prediction, which sometimes leads to accuracy degradation [pmlr-v80-kumar18a, karandikar2021soft, NEURIPS2022_33d6e648, marx2023calibration, Kuleshov2015]. To theoretically guarantee the calibration, our analysis takes a different perspective. We provide guarantees for the smooth CE without relying on such a post-processing approach similar to recalibration. This is achieved by combining a uniform convergence bound and a functional gradient characterization of the training smooth CE. Then, our analysis simultaneously guarantees both the smooth CE and accuracy for several practical algorithms.
Prior work has connected the post-processing gap to the population smooth CE [blasiok2023unify, blasiok2023when], making it less applicable to algorithms trained on finite datasets. Similarly, although gruber2022better discussed post-processing, their approach faces the same limitation. In contrast, our analysis directly targets the smooth CE from finite data. From a generalization perspective, NEURIPS2024_futami developed algorithm-dependent bounds for binning ECE using information-theoretic techniques. By contrast, we derive a uniform convergence bound for smooth CE, which applies more broadly beyond binning. Moreover, their generalization bound for ECE converges at a rate $\mathcal{O}(\log n/n^{1/3})$, which is slower than our rate. We conjecture that this slower convergence arises from the nonparametric estimation of conditional probabilities via binning, whereas smooth CE avoids such estimation and hyperparameter dependence. Finally, their results focus only on the generalization gap and do not address when the training ECE itself becomes small.
Although in recent studies, new boosting algorithms have been proposed to improve various notions of calibration [johnson18a, pmlr-v202-globus-harris23a], our analysis provides a theoretical explanation for why existing gradient boosting already yields strong calibration performance. Unlike prior works on boosting, which primarily focus on accuracy on the basis of functional gradients [Zhang2005BoostingWE, nitanda18a, nitanda2019gradient], we leverage functional gradients to analyze calibration, offering a novel perspective on the behavior of these algorithms in calibration. While our analysis assumes a constant stepsize, the choice of stepsize in boosting is crucial for achieving better performance [telgarsky2013margins]. Extending our calibration analysis beyond the constant stepsize setting is an important direction for future work.
In several works, the post-processing has been highlighted in achieving advanced calibration objectives, such as multicalibration [johnson18a, hansen2024multicalibration, blasiok2023loss, pmlr-v202-globus-harris23a]. Whereas our analysis focuses on smooth CE, this metric is closely related to these variants of CE from the viewpoint of low-degree calibration [gopalan2022low]. We believe our findings may offer new insights into these metrics, which we leave for future exploration.
6 Conclusion and limitation
This work presented the first rigorous guarantees on the smooth CE for several widely used learning algorithms. Our analysis proceeds in two stages: we first derive a uniform convergence bound for the smooth CE, and then upper-bound the training smooth CE via functional gradients. We demonstrate that algorithms that can control the functional gradient simultaneously achieve a small smooth CE and misclassification rate. Despite these contributions, several limitations remain. First, our bounds are derived using a uniform bound over the post-processing function $h$, not the Lipschitz function class, which may result in loose estimates. More refined analyses that better align with practical performance would be valuable. Moreover, our analysis relies on a strong margin assumption, which requires a well-separated data distribution. Since calibration is a rather weak notion compared with accuracy, extending the analysis to more realistic and weaker conditions is an important direction for future work. Finally, our analysis is limited to binary classification as the smooth CE is designed for this setting. Extending our analysis to the multiclass setting should be explored in future work.
Acknowledgements
FF was supported by JSPS KAKENHI Grant Number JP23K16948. FF was supported by JST, PRESTO Grant Number JPMJPR22C8, Japan. This research is supported by the National Research Foundation, Singapore, Infocomm Media Development Authority under its Trust Tech Funding Initiative, and the Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-004). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore, Infocomm Media Development Authority, and the Ministry of Digital Development and Information.
Appendix A Additional facts about proper losses
Here, we describe the general form of a proper loss function and its relation to the post-processing gap. Given a proper loss function $\ell$, it can always be represented in terms of a convex function as follows:
$\ell(p,y)=-\phi(p)-\nabla\phi(p)\cdot(y-p),$ (29)
where $\phi:[0,1]\to\mathbb{R}$ is a convex function and $\nabla\phi(p)$ denotes a subgradient at $p$. This representation is known as the Savage representation [Gneiting2007]. blasiok2023when further showed that, in the binary case, the subgradient satisfies $\nabla\phi(p)=\ell(p,0)-\ell(p,1)$, and referred to $\nabla\phi(p)$ as the dual of $p$ ($\text{dual}(p)$). Hereinafter, we assume that $\phi$ is differentiable, since the convex functions corresponding to commonly used losses such as the squared loss and the log loss are differentiable.
Table 1: Savage representations for the cross-entropy Loss and squared loss
cross-entropy Loss Squared Loss
$\ell(p,y)$ $y\ln p-(1-y)\ln(1-p)$ $(y-p)^{2}$
$\phi(v)$ $p\ln p+(1-p)\ln(1-p)$ p(p1)p(p-1)
$\text{dual}(p)$ $\ln\left(\frac{p}{1-p}\right)$ 2p12p-1
$\psi(s)$ $\ln(1+e^{s})$ $\begin{cases}0,&s<-1\\ \frac{(s+1)^{2}}{4},&-1\leq s\leq 1\\ s,&s>1\end{cases}$
$\ell^{\psi}(s,y)$ $\ln(1+e^{s})-y\cdot s$ $\begin{cases}-ys,&s<-1\\ (y-(s+1)/2)^{2},&-1\leq s\leq 1\\ (1-y)s,&t>1\end{cases}$
$\mathrm{pred}_{\psi}(s)(=\nabla\psi(s))$ $\frac{e^{s}}{1+e^{s}}$ $\begin{cases}0,&s<-1\\ \frac{s+1}{2},&-1\leq s\leq 1\\ 1,&s>1\end{cases}$
We define the convex conjugate of a function $\phi(p)$ as follows: for all $s\in\mathbb{R}$,
$\psi(s)=\sup_{p\in[0,1]}\left\{s\cdot p-\phi(p)\right\},$ (30)
where $\psi$ is a convex function. We refer to $s$ as a score, as will be explained later. Using this notation, we define the dual loss $\ell^{\psi}:\mathbb{R}\times\mathcal{Y}\to\mathbb{R}$ as
$\ell^{\psi}(s,y)\coloneqq\psi(s)-s\cdot y.$ (31)
The score $s$ is linked to the probability $p$ via the dual mapping $\text{dual}(p)=\nabla\phi(p)=\ell(p,0)-\ell(p,1)$. By Fenchel–Young duality, this relationship is inverted as $p=\nabla\psi(s)$, which we also denote as $p=\text{pred}_{\psi}(s)$.
With these definitions, the proper loss can be equivalently written as
$\ell(p,y)=\ell^{\psi}(\text{dual}(p),y)=\psi(\text{dual}(p))-\text{dual}(p)\cdot y.$ (32)
For details and proofs, see blasiok2023when. We summarize the key properties of these expressions in Table 1.
Similarly, we can define the dual post-processing gap for the proper scoring rule [blasiok2023when]:
Definition 3 (Dual-post processing gap).
Assume that $\Psi$ is a differentiable and convex function with derivative $\nabla\psi(t)\in[0,1]$ for all $t\in\real$. Assume that $\psi$ is $\lambda$ smooth function. Given $\psi$, $\ell^{\psi}$, $g:\mathcal{X}\to\real$, and $\mathcal{D}$, we define the dual post-processing gap as
$\mathrm{pGap}^{(\psi,\lambda)}(g,\mathcal{D})\coloneqq\mathbb{E}[\ell^{\psi}(g(X),Y)]-\inf_{h\in\mathrm{Lip}_{1}(\real,[-1/\lambda,1/\lambda])}\mathbb{E}[\ell^{\psi}(g(X)+h(g(X)),Y)]$ (33)
When considering the cross-entropy loss, this dual post-processing gap corresponds to improving the logit function.
Definition 4 (Dual smooth calibration).
Consider the same setting as the definition of the dual-post processing gap. Given $\psi$ and $g$, we define $f(\cdot)=\nabla\psi(g(\cdot))$. We then define the dual calibration error of $g$ as
$\mathrm{smCE}^{(\psi,\lambda)}(g,\mathcal{D})\coloneqq\sup_{h\in\mathrm{Lip}_{L=\lambda}(\real,[-1,1])}\mathbb{E}\eta(g(X))\cdot(Y-f(X))$ (34)
Then, similarly to the relationship between the smooth ECE and the post-processing gap, the following holds: if $\psi$ is a $\lambda$-smooth function, then we have
$\mathrm{smCE}^{(\psi,\lambda)}(g,\mathcal{D})^{2}/2\leq\lambda\mathrm{pGap}^{(\psi,\lambda)}(g,\mathcal{D})^{2}/2\leq\mathrm{smCE}^{(\psi,\lambda)}(g,\mathcal{D})$ (35)
and
$\mathrm{smCE}(f,\mathcal{D})\leq\mathrm{smCE}^{(\psi,\lambda)}(g,\mathcal{D})$ (36)
holds.
In the case of the cross-entropy loss, we have $\psi(s)=\log(1+e^{s})$, which is $1/4$-smooth. For the squared loss, $\psi$ is $1/2$-smooth. Therefore, the dual smooth calibration and the smooth calibration are essentially equivalent in both cases.
Appendix B Discussion about Calibration metrics
Here, we introduce several definitions of calibration metrics and their relationship to the smooth CE, which is the primary focus of our paper.
B.1 True distance calibration and smooth CE
Given a distribution $\mathcal{D}$ and predictor $f$, we consider the distribution induced by $f$, which is the joint distribution of prediction-label pairs $(f(x),y)\in[0,1]\times\{0,1\}$, denoted by $\mathcal{D}_{f}$.
Definition 5 (Perfect calibration [blasiok2023when]).
We say that a prediction-label distribution $\Gamma$ over $0,1]\times\{0,1\}$ is perfectly calibrated if $\mathbb{E}_{(v,y)\sim\Gamma}[y|v]=v$. Moreover, given $\mathcal{D}$ over $\mathcal{X}\times\{0,1\}$, we say that $f$ is perfectly calibrated with respect to $\mathcal{D}$ if $\mathcal{D}_{f}$ is perfectly calibrated.
We denote the set of perfectly calibrated predictors with respect to $\mathcal{D}$ by $\mathrm{cal}(\mathcal{D})$.
Next, we introduce the true calibration distance, defined as the distance to the closest calibrated predictor:
Definition 6 (True calibration distance).
The true distance of a predictor $f$ from calibration is defined as
$\mathrm{dCE}_{\mathcal{D}}(f):=\inf_{g\in\mathrm{cal}(\mathcal{D})}\mathbb{E}_{\mathcal{D}}|f(x)-g(x)|.$ (37)
This is the $l_{1}$ metric, which possesses many desirable properties for measuring calibration; see blasiok2023unify for details. However, it is not practically usable because $\mathrm{cal}(\mathcal{D})$ may be non-convex, and the metric depends on the domain $\mathcal{X}$, whereas calibration metrics typically depend only on $\mu(\mathcal{D},f)$.
blasiok2023unify proposed that any calibration metric should satisfy the following three criteria:
(i) Prediction-only access
A calibration metric should depend only on the distribution over $(f(x),y)\sim\mathcal{D}_{f}$, and not on the distribution over $\mathcal{X}$. If it meets this requirement, we say that a calibration metric $\mu$ satisfies the Prediction-only access (PA) property. Note that the true calibration distance depends on $(x,f(x),y)$ and is therefore called the sample access (SA) model.
(ii) Consistency
Definition 7.
For $c>0$, a calibration metric $\mu$ is said to satisfy $c$-robust completeness if there exists a constant $a\geq 0$ such that for every distributionDD on $\mathcal{X}\times\{0,1\}$, and predictor $f\in\mathcal{F}$
$\mu(\mathcal{D},f)\leq a\left(\mathrm{dCE}_{\mathcal{D}}(f)\right)^{c}.$ (38)
Definition 8.
For $s>0$, a calibration metric $\mu$ satisfies $s$-robust soundness if there exists a constant $b\geq 0$ such that for every distribution $D$ on $\mathcal{X}\times\{0,1\}$, and predictor $f\in\mathcal{F}$,
$\mu(\mathcal{D},f)\geq b\left(\mathrm{dCE}_{\mathcal{D}}(f)\right)^{s}.$ (39)
Definition 9.
A calibration metric $\mu$ is said to be $(c,s)$-consistent if it satisfies both $c$-robust completeness and $s$-robust soundness.
Thus, consistent calibration metrics are polynomially related to the true distance from calibration. blasiok2023unify showed that any PA model must satisfy $s/c\geq 2$.
(iii) Efficiency
A calibration metric $\mu$ is said to be efficient if it can be computed to accuracy $\epsilon$ in time $\text{poly}(1/\epsilon)$ using $\text{poly}(1/\epsilon)$ random samples from $\mathcal{D}_{f}$.
blasiok2023unify showed that the smooth CE satisfies all of these properties. In particular, with respect to consistency, the smooth CE satisfies
$\mathrm{smCE}(f,\mathcal{D})\leq\mathrm{dCE}_{\mathcal{D}}(f)\leq 4\sqrt{2\mathrm{smCE}(f,\mathcal{D})}$ (40)
$\Leftrightarrow\frac{1}{32}\mathrm{dCE}_{\mathcal{D}}(f)^{2}\leq\mathrm{smCE}(f,\mathcal{D})\leq\mathrm{dCE}_{\mathcal{D}}(f)$ (41)
which means that the smooth CE is $(c=1,s=2)$-consistent and achieves $s/c=2$.
B.2 ECE and binning estimator
Recall the definition of the expected calibration error (ECE):
$\mathrm{ECE}_{\mathcal{D}}(f):=\mathbb{E}\left[\left|\mathbb{E}[Y|f(X)]-f(X)\right|\right].$ (42)
The binning estimator is commonly used to approximate this. Given a partition $\mathcal{I}=\{I_{1},\dots,I_{m}\}$ of $[0,1]$ into intervals, we define:
$\mathrm{binnedECE}_{\mathcal{D}}(\Gamma,\mathcal{I})=\sum_{j\in[m]}|\mathbb{E}_{(V,Y\sim\Gamma)}[(V-Y)\mathbbm{1}(V\in I_{j})]|.$ (43)
It has been shown in Lemma 4.7 of blasiok2023unify that
$\mathrm{dCE}_{\mathcal{D}}(f)\leq\mathrm{ECE}_{\mathcal{D}}(f)$ (44)
demonstrating robust soundness. However, ECE does not satisfy robust completeness; blasiok2023unify proved (Lemma 4.8) that for any $\epsilon\in\real^{+}$, there exists a distribution $\mathcal{D}$ such that $\mathrm{dCE}_{\mathcal{D}}(f)\leq\epsilon$ but $\mathrm{ECE}_{\mathcal{D}}(f)\geq 1/2-\epsilon$. This also highlights the discontinuity of ECE, which hinders its estimation from finite samples.
blasiok2023unify showed that by accounting for bin widths, the binning ECE satisfies consistency:
$\text{intCE}(f)\coloneqq\min_{\mathcal{I}}(\mathrm{binnedECE}_{\mathcal{D}}(\Gamma,\mathcal{I})+w_{\Gamma}(\mathcal{I}))$ (45)
where
$w_{\Gamma}(\mathcal{I})\coloneqq\sum_{j\in[m]}|\mathbb{E}_{(V,Y\sim\Gamma)}w(I_{j})\mathbbm{1}(V\in I_{j})|$ (46)
and $w(I)$ denotes the width of interval $I$.
Then, the following holds (Theorem 6.3 in blasiok2023unify):
$\mathrm{dCE}_{\mathcal{D}}(f)\leq\text{intCE}(f)\leq 4\sqrt{\mathrm{dCE}_{\mathcal{D}}(f)}$ (47)
Thus, intCE satisfies $(1/2,1)$-consistency. blasiok2023unify also provided an estimator for $\text{intCE}(\Gamma)$.
As we have seen, bounding the smooth CE leads to a bound on $\mathrm{dCE}_{\mathcal{D}}(f)$, which in turn bounds $\text{intCE}(f)$—an optimized estimator for the binned ECE.
Finally, we remark that since $\mathbb{E}[Y|f=v]$ is continuous, we can relate the ECE and the smooth CE as follows:
Theorem 5.
Suppose that $\mathbb{E}[Y|f=v]$ satisfies $L$-Lipschitz continuity. Then the following relation holds:
$\frac{1}{2}\mathrm{smCE}(f,\mathcal{D})\leq\mathrm{ECE}_{\mathcal{D}}(f)\leq(1+2\sqrt{2(1+L)})\sqrt{2\mathrm{smCE}(f,\mathcal{D})}.$ (48)
Thus, controlling the smooth CE also leads to controlling the ECE when the underlying conditional probability function is continuous.
Proof.
We first prove the following inequality:
$\mathrm{ECE}_{\mathcal{D}}(f)\leq\mathrm{binnedECE}_{\mathcal{D}}(\Gamma,\mathcal{I})+(1+L)w_{\Gamma}(\mathcal{I}).$ (49)
This can be derived using Lemma 4 in NEURIPS2024_futami. For completeness, we outline the key steps of the analysis below:
$\mathrm{ECE}_{\mathcal{D}}(f)=\mathbb{E}\left[|\mathbb{E}[Y|f=V]-V|\right]$ $=\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|f=V]-V||V\in I_{i}]$
$=\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|f=V]-\mathbb{E}[V|V\in I_{i}]$
$\quad+\mathbb{E}[V|V\in I_{i}]-V||V\in I_{i}]$
$\leq\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V]-\mathbb{E}[Y|V\in I_{i}]||V\in I_{i}]$
$\quad+\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V\in I_{i}]-\mathbb{E}[V|V\in I_{i}]||V\in I_{i}]$
$\quad+\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[V|V\in I_{i}]-V||V\in I_{i}].$ (50)
As for the second term, it is evident that
$\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}|\mathbb{E}[Y|V\in I_{i}]-\mathbb{E}[V|V\in I_{i}]|=\mathrm{binnedECE}_{\mathcal{D}}(\Gamma,\mathcal{I})$ (51)
and as for the third term,
$\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[V|V\in I_{i}]-V||V\in I_{i}]\leq\sum_{i=1}^{m}\mathbb{E}[w(I_{i})\mathbbm{1}_{V\in I_{i}}]=w(\Gamma).$ (52)
Finally, the first term can be bounded by using the Lipschitz continuity,
$\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V]-\mathbb{E}[Y|V\in I_{i}]||V\in I_{i}]$ (53)
$\leq\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V]-\mathbb{E}[Y|V_{\mathcal{I}}]||V\in I_{i}]$ (54)
$\quad+\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|\mathbb{E}[Y|V_{\mathcal{I}}]-\mathbb{E}[Y|V\in I_{i}]||V\in I_{i}]$ (55)
$\leq L\sum_{i=1}^{m}P(V\in I_{i})\mathbb{E}[|V-\mathbb{E}[V|V\in I_{i}]||V\in I_{i}]\leq Lw(\Gamma)$ (56)
where
$V_{\mathcal{I}}\coloneqq\sum_{i=1}^{m}V_{I_{i}}\cdot\mathbbm{1}_{V\in I_{i}}=\sum_{i=1}^{m}\mathbb{E}[V|V\in I_{i}]\cdot\mathbbm{1}_{V\in I_{i}}.$ (57)
The term $\mathbb{E}[|\mathbb{E}[Y|V_{\mathcal{I}}]-\mathbb{E}[Y|V\in I_{i}]||V\in I_{i}]=0$ holds by the definition of conditional expectation. Thus, we have established Eq. equation 49.
Next, we consider taking the infimum over all partitions $\mathcal{I}$. By combining Claim 6.6 and Lemma 6.7 in blasiok2023unify and Eq. equation 49, we obtain:
[…]
We remark that if the underlying conditional distribution satisfies a continuity condition, then the ECE becomes a consistent calibration metric.
B.3 Other metrics
As discussed in blasiok2023unify and gopalan2022low, the smooth CE is a special case of the weighted calibration error introduced by gopalan2022low.
Definition 10 (Weighted Calibration Error).
Let $\mathcal{M}$ be a class of functions $h:[0,1]\to\mathbb{R}$. The calibration error relative to $\mathcal{M}$ is defined as
$\mathrm{CE}_{\mathcal{D}}(f,\mathcal{M})\coloneqq\sup_{h\in\mathcal{M}}\mathbb{E}\left[h(f(X))\cdot(Y-f(X))\right]=\sup_{h\in\mathcal{M}}\left\langle h(f(X)),Y-f(X)\right\rangle_{L_{2}(\mathcal{D})}.$ (59)
In this view, the smooth CE corresponds to $\mathrm{smCE}(f,\mathcal{D})=\mathrm{CE}_{\mathcal{D}}(f,\mathrm{Lip}_{1}([0,1],[-1,1]))$.
For the dual smooth CE, the inverse of the sigmoid function is $\sigma^{-1}(x)=\log(x/(1-x))$. When $g(x)=\sigma^{-1}(f(x))$, we have:
$\mathrm{smCE}^{\sigma}(g,\mathcal{D})=\mathrm{CE}_{\mathcal{D}}(f,\mathrm{Lip}_{1/4}(\real,[-1,1])\circ\sigma^{-1})$
Another important function class is the RKHS $\mathcal{H}$ associated with a positive definite kernel $k$. This space is equipped with the feature map $\phi:\mathbb{R}\to\mathcal{H}$ satisfying $\langle h,\phi(v)\rangle_{\mathcal{H}}=h(v)$. The associated kernel $k:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is defined by $k(u,v)=\langle\phi(u),\phi(v)\rangle_{\mathcal{H}}$.
Let $\mathcal{H}_{1}\coloneqq\{h\in\mathcal{H}|\|h\|_{\mathcal{H}}\leq 1\}$. We define the kernel calibration error as $\mathrm{CE}_{\mathcal{D}}(f,\mathcal{H}_{1})$.
Given samples $\{(v_{1},y_{1}),\dots,(v_{n},y_{n}))\}$ where $v_{i}=f(x_{i})$, the kernel CE under finite samples is defined as
$\hat{\mathrm{CE}}_{\mathcal{D}}(f,\mathcal{W}_{H})^{2}\coloneqq\frac{1}{n^{2}}\sum_{i,j}(y_{i}-v_{i})k(v_{i},v_{j})(y_{j}-v_{j}).$ (60)
This was first proposed by pmlr-v80-kumar18a as the maximum mean calibration error (MMCE).
A key kernel is the Laplace (exponential) kernel $k_{\mathrm{Lap}}(u,v)=\exp(-|u-v|)$, for which it has been shown that
$\frac{1}{3}\mathrm{smCE}(f,\mathcal{D})\leq\mathrm{CE}_{\mathcal{D}}(f,\mathcal{H}_{1}).$ (61)
On the other hand, the Gaussian kernel does not upper bound the smooth CE and has several limitations; see blasiok2023unify for details.
We numerically evaluate the behavior of MMCE with the Laplace kernel and smooth CE in Appendix I.
Appendix C Proof of Section 3
C.1 Proof of Theorem 1
Before the formal proof, we provide a proof sketch to highlight the key differences from standard generalization bounds.
First, we reformulate the smooth CE as a linear convex optimization [hu2024testing]: let $v_{i}=f(x_{i})$ and $\omega=(\omega_{1},\dots,\omega_{n})\in\mathbb{R}^{n}$. Then $\mathrm{smCE}(f,S)=\max_{\omega}\sum_{i=1}^{n}(y_{i}-v_{i})\omega_{i}/n$ subject to the constraints $|\omega_{i}|\leq 1$, $\forall i$, and $|\omega_{i}-\omega_{j}|\leq|v_{i}-v_{j}|$, $\forall i,j$. Let $\omega^{*}$ denote one such solution, and we define $\phi(f,z_{i})\coloneqq(y_{i}-v_{i})\omega_{i}^{*}$. Then, we express $\phi(f,S)\!=\!\frac{1}{n}\sum_{i=1}^{n}\phi(f,z_{i})$.
Although our analysis follows the structure of classical generalization bounds, the dependence among $\{\phi(f,z_{i})\}_{i=1}^{n}$ induced by $\omega^{*}$ through the optimization precludes the use of standard concentration inequalities based on independence. Instead of Hoeffding’s inequality, we use the bounded difference inequality to show $\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|\leq\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}}\sim D^{2n}}\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|+\sqrt{\log\delta^{-1}/n}$. This requires studying the stability of the above convex problem.
We would like to apply a symmetrization argument using i.i.d. Rademacher variables $\sigma_{i}\in{\pm 1}$ and evaluate $\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}}\!\sim D^{2n}}\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\left[\phi(f,Z^{\prime}_{i})\!-\phi(f,Z_{i})\right]$. However, this technique is not directly applicable in our setting due to dependencies among the terms $\{\phi(f,z_{i})\}_{i=1}^{n}$.
To address this, we introduce $\{\sigma_{i}\}$s in a way that preserves the structure of the linear convex formulation of smooth CE. Since the resulting bound does not take the form of standard Rademacher complexity, we discretize the hypothesis class $\mathcal{F}$ and derive a covering number bound.
Proof.
We begin by leveraging the reformulation of the smooth CE as a linear program introduced by blasiok2023unify. Given a dataset $S=\{(x_{i},y_{i})\}_{i=1}^{n}$ and defining $v_{i}=f(x_{i})$, it is known that the smooth CE on $S$ corresponds to the optimal value of the following optimization problem:
$\mathrm{smCE}(f,S)=\max_{\{\omega_{i}\}}\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\omega_{i}$ (62)
subject to the constraints:
$1\leq\omega_{i}\leq 1,\quad\forall i,\quad|\omega_{i}-\omega_{j}|\leq|v_{i}-v_{j}|,\quad\forall i,j.$
This is a convex linear program whose optimal value exists; see Lemma 7.6 in blasiok2023unify. Note that the optimal solution is not necessarily unique. Nevertheless, given a function $f$ and dataset $S$, the value $\mathrm{smCE}(f,S)$ is uniquely determined. Let $\omega^{*}_{i}$ denote one such optimal solution. Then, we can write:
$\mathrm{smCE}(f,S)=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\omega^{*}_{i}=\frac{1}{n}\sum_{i=1}^{n}\phi(f,z_{i})$ (63)
where $z_{i}=(x_{i},y_{i})$. We denote $\mathrm{smCE}(f,S_{\mathrm{tr}})=\phi(f,S_{\mathrm{tr}})$ and $\mathrm{smCE}(f,S_{\mathrm{te}})=\phi(f,S_{\mathrm{te}})$.
As discussed above, although the optimizer $\omega^{*}$ may not be unique, the value $\mathrm{smCE}(f,S)$ is uniquely determined for a fixed $f$ and dataset $S$. Therefore, the uniform upper bound on the smooth CE can be controlled via the supremum over $f$ when the training and test datasets are fixed. We thus focus on the following inequality:
$|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|\leq\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|$ (64)
Following the standard approach in uniform convergence analysis, we derive the convergence in expectation as follows:
Lemma 1.
Under the same setting as Theorem 1, with probability $1-\delta$, we have
$\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|\leq\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}}\sim D^{2n}}\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|+2\sqrt{\frac{\log\frac{1}{\delta}}{n}}.$ (65)
The proof of this lemma is provided in Appedix C.2.
Note that this is not a consequence of Hoeffding’s inequality, since
$\phi(f,S_{\mathrm{tr}})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\omega^{*}_{i}=\frac{1}{n}\sum_{i=1}^{n}\phi(f,z_{i})$ (66)
and thus the terms $\phi(f,z_{i})$ are dependent through $\omega^{*}_{i}$, which is a solution of the linear program in Eq. equation 62. Since there are $n$ dependent variables $\phi(f,z_{i})$, Hoeffding’s inequality, which requires independence, is not applicable. To establish Lemma 1, we instead employ McDiarmid’s inequality combined with a bounded difference argument; see the proof in Appendix C.2.
Our proof proceeds in three steps: (1) we introduce Rademacher random variables, (2) we evaluate the exponential moment to control the empirical process induced by these variables, and (3) we refine the exponential moment bound via a chaining argument.
We introduce Rademacher random variables to control the empirical process in Lemma 1. Following standard generalization analysis, we aim to apply the symmetrization technique. However, the standard formulation
$\mathbb{E}_{(S_{n},S^{\prime}_{n})\sim D^{2n}}\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\left[\sigma_{i}\phi(f,Z^{\prime}_{i})-\sigma_{i}\phi(f,Z_{i})\right]$ (67)
is not suitable for Step (2), as it complicates the evaluation of the exponential moment.
To understand this, let us explicitly express the uniform generalization gap as follows:
$\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}}\sim D^{2n}}\sup_{f\in\mathcal{F}}\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})$ (68)
$=\mathbb{E}_{(S_{n},S^{\prime}_{n})\sim D^{2n}}\sup_{f\in\mathcal{F}}\left(\max_{\omega}\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f(x_{i}))\omega_{i}-\max_{\omega^{\prime}}\frac{1}{n}\sum_{i=1}^{n}(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{\prime}_{i}\right)$ (69)
$=\mathbb{E}_{(S_{n},S^{\prime}_{n})\sim D^{2n}}\sup_{f\in\mathcal{F}}\max_{\omega}\min_{\omega^{\prime}}\frac{1}{n}\sum_{i=1}^{n}\left[(y_{i}-f(x_{i}))\omega_{i}-(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{\prime}_{i}\right]$ (70)
Here, we used the reformulation of the smooth CE over the training dataset via a Lipschitz function. For simplicity, we omit the constraints of the linear program. We now introduce Rademacher variables and demonstrate the difficulty in evaluating the exponential moment.
To simplify the discussion, let us consider the case $n=2$. For a fixed dataset $S$ and function $f$, Eq. equation 67 under the expression of Eq. equation 68 is written as:
$\mathbb{E}_{\sigma}\max_{\omega}\min_{\omega^{\prime}}\frac{\sigma_{1}}{2}\left[(y_{1}-f(x_{1}))\omega_{1}-(y^{\prime}_{1}-f(x^{\prime}_{1}))\omega^{\prime}_{1}\right]+\frac{\sigma_{2}}{2}\left[(y_{2}-f(x_{2}))\omega_{2}-(y^{\prime}_{2}-f(x^{\prime}_{2}))\omega^{\prime}_{2}\right]$ (71)
$=\frac{1}{8}\max_{\omega}\min_{\omega^{\prime}}(y_{1}-f(x_{1}))\omega_{1}+(y_{2}-f(x_{2}))\omega_{2}-(y^{\prime}_{1}-f(x^{\prime}_{1}))\omega^{\prime}_{1}-(y^{\prime}_{2}-f(x^{\prime}_{2}))\omega^{\prime}_{2}$ (72)
$+\frac{1}{8}\max_{\omega}\min_{\omega^{\prime}}(y_{1}-f(x_{1}))\omega_{1}-(y_{2}-f(x_{2}))\omega_{2}-(y^{\prime}_{1}-f(x^{\prime}_{1}))\omega^{\prime}_{1}+(y^{\prime}_{2}-f(x^{\prime}_{2}))\omega^{\prime}_{2}$ (73)
$+\frac{1}{8}\max_{\omega}\min_{\omega^{\prime}}-(y_{1}-f(x_{1}))\omega_{1}+(y_{2}-f(x_{2}))\omega_{2}+(y^{\prime}_{1}-f(x^{\prime}_{1}))\omega^{\prime}_{1}-(y^{\prime}_{2}-f(x^{\prime}_{2}))\omega^{\prime}_{2}$ (74)
$+\frac{1}{8}\max_{\omega}\min_{\omega^{\prime}}-(y_{1}-f(x_{1}))\omega_{1}-(y_{2}-f(x_{2}))\omega_{2}+(y^{\prime}_{1}-f(x^{\prime}_{1}))\omega^{\prime}_{1}+(y^{\prime}_{2}-f(x^{\prime}_{2}))\omega^{\prime}_{2}$ (75)
$\neq 0$ (76)
This non-zero result indicates the challenge in applying exponential moment analysis for Step (2). In standard analysis, this type of expectation is typically zero, which enables the use of tools such as Massart’s lemma. Hence, the standard symmetrization technique fails in this setting.
Therefore, we must develop an alternative symmetrization strategy. To illustrate the idea, we begin by introducing only $\sigma_{1}$ and consider the following expression:
Eq. equation 70 (77)
$=\mathbb{E}_{(S_{n},S^{\prime}_{n})\sim D^{2n}}\mathbb{E}_{\sigma_{1}}\sup_{f\in\mathcal{F}}\max_{\omega}\min_{\omega^{\prime}}\Bigg[$ (78)
$\quad\frac{\sigma_{1}}{n}\left((y_{1}-f(x_{1}))\left(\frac{1+\sigma_{1}}{2}\omega_{1}+\frac{1-\sigma_{1}}{2}\omega^{\prime}_{1}\right)-(y^{\prime}_{1}-f(x^{\prime}_{1}))\left(\frac{1+\sigma_{1}}{2}\omega^{\prime}_{1}+\frac{1-\sigma_{1}}{2}\omega_{1}\right)\right)$ (79)
$\quad+\frac{1}{n}\sum_{i=2}^{n}\left((y_{i}-f(x_{i}))\omega_{i}-(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{\prime}_{i}\right)\Bigg]$ (80)
Here, $\sigma_{1}$ is a Rademacher random variable. This equality holds because, when $\sigma_{1}=1$, the maximization and minimization are unchanged; when $\sigma_{1}=-1$, the roles of $\omega_{1}$ and $\omega^{\prime}_{1}$ are exchanged. However, since we are averaging over $S$ and $S^{\prime}$, and these datasets are i.i.d., such swapping does not affect the distribution of the expectation. Therefore, the equality holds. This method of introducing Rademacher variables differs fundamentally from the standard symmetrization approach.
We then introduce i.i.d. Rademacher random variables $\{\sigma_{i}\}_{i=1}^{n}$ and consider the following expression:
Eq. equation 70 (81)
$=\mathbb{E}_{(S_{n},S^{\prime}_{n})\sim D^{2n}}\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\max_{\omega}\min_{\omega^{\prime}}$ (82)
$\quad\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\left[(y_{i}-f(x_{i}))\left(\frac{1+\sigma_{i}}{2}\omega_{i}+\frac{1-\sigma_{i}}{2}\omega^{\prime}_{i}\right)-(y^{\prime}_{i}-f(x^{\prime}_{i}))\left(\frac{1+\sigma_{i}}{2}\omega^{\prime}_{i}+\frac{1-\sigma_{i}}{2}\omega_{i}\right)\right]$ (83)
An important property is that for fixed $f$ and dataset $S$, we have:
$\mathbb{E}_{\sigma}\max_{\omega}\min_{\omega^{\prime}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\left[(y_{i}-f(x_{i}))\left(\frac{1+\sigma_{i}}{2}\omega_{i}+\frac{1-\sigma_{i}}{2}\omega^{\prime}_{i}\right)\right$. (84)
$\qquad\left.-(y^{\prime}_{i}-f(x^{\prime}_{i}))\left(\frac{1+\sigma_{i}}{2}\omega^{\prime}_{i}+\frac{1-\sigma_{i}}{2}\omega_{i}\right)\right]=0.$ (85)
This cancellation becomes evident in the case $n=2$: by expanding the left-hand side above, we obtain
$\frac{1}{8}\max_{\omega}\min_{\omega^{\prime}}(y_{1}-f(x_{1}))\omega_{1}+(y_{2}-f(x_{2}))\omega_{2}-(y^{\prime}_{1}-f(x^{\prime}_{1}))\omega^{\prime}_{1}-(y^{\prime}_{2}-f(x^{\prime}_{2}))\omega^{\prime}_{2}$ (86)
$+\frac{1}{8}\max_{\omega}\min_{\omega^{\prime}}(y_{1}-f(x_{1}))\omega_{1}+(y^{\prime}_{2}-f(x^{\prime}_{2}))\omega_{2}-(y^{\prime}_{1}-f(x^{\prime}_{1}))\omega^{\prime}_{1}-(y_{2}-f(x_{2}))\omega^{\prime}_{2}$ (87)
$+\frac{1}{8}\max_{\omega}\min_{\omega^{\prime}}(y^{\prime}_{1}-f(x^{\prime}_{1}))\omega_{1}+(y_{2}-f(x_{2}))\omega_{2}-(y_{1}-f(x_{1}))\omega^{\prime}_{1}-(y^{\prime}_{2}-f(x^{\prime}_{2}))\omega^{\prime}_{2}$ (88)
$+\frac{1}{8}\max_{\omega}\min_{\omega^{\prime}}(y^{\prime}_{1}-f(x^{\prime}_{1}))\omega_{1}+(y^{\prime}_{2}-f(x^{\prime}_{2}))\omega_{2}-(y_{1}-f(x_{1}))\omega^{\prime}_{1}-(y_{2}-f(x_{2}))\omega^{\prime}_{2}$ (89)
$=0.$ (90)
This is because the structure of the linear convex problem ensures symmetric cancellation when Rademacher variables are introduced. Therefore, we can proceed the analysis of the exponential moment of Step (2).
We define the integrated empirical process as
$R(\mathcal{F},S_{n},S^{\prime}_{n})\coloneqq\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\max_{\omega}\min_{\omega^{\prime}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,Z_{i},Z^{\prime}_{i}),$ (91)
where
$\phi(f,Z_{i},Z^{\prime}_{i})\coloneqq(y_{i}-f(x_{i}))\left(\frac{1+\sigma_{i}}{2}\omega_{i}+\frac{1-\sigma_{i}}{2}\omega^{\prime}_{i}\right)-(y^{\prime}_{i}-f(x^{\prime}_{i}))\left(\frac{1+\sigma_{i}}{2}\omega^{\prime}_{i}+\frac{1-\sigma_{i}}{2}\omega_{i}\right).$ (92)
We now derive a bound using a variant of the Massart lemma. For simplicity, assume the function class $\mathcal{F}$ has finite cardinality $|\mathcal{F}|$. This will later be replaced by a covering number. Then for all $\lambda>0$, we have:
$R(\mathcal{F},S_{n},S^{\prime}_{n})$ $=\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\max_{\omega}\min_{\omega^{\prime}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,Z_{i},Z^{\prime}_{i})$ (93)
$\leq\mathbb{E}_{\sigma}\frac{1}{\lambda}\log\sum_{f\in\mathcal{F}}\exp\left(\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,Z_{i},Z^{\prime}_{i})\right)$ (94)
$\leq\frac{1}{\lambda}\log\sum_{f\in\mathcal{F}}\mathbb{E}_{\sigma}\exp\left(\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,Z_{i},Z^{\prime}_{i})\right)$ (95)
$\leq\frac{1}{\lambda}\log\left(|\mathcal{F}|\cdot\mathbb{E}_{\sigma}\exp\left(\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,Z_{i},Z^{\prime}_{i})\right)\right).$ (96)
Since the expectation of the exponential term is 0 from Eq. equation 84, we apply McDiarmid’s inequality:
Lemma 2.
Under the above setting, we have
$\mathbb{E}_{\sigma}\exp\left(\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,Z_{i},Z^{\prime}_{i})\right)\leq\exp\left(\frac{\lambda^{2}}{2n}\right).$ (97)
The proof of this theorem is provided in Appendix C.3. Using this, we obtain
$R(\mathcal{F},S_{n},S^{\prime}_{n})\leq\frac{\log|\mathcal{F}|}{\lambda}+\frac{\lambda}{2n}.$ (98)
Optimizing over $\lambda$, we find
$R(\mathcal{F},S_{n},S^{\prime}_{n})\leq\sqrt{\frac{2\log|\mathcal{F}|}{n}}.$ (99)
This is a modified version of Massart’s lemma since the left-hand side does not represent the classical Rademacher complexity.
Since $\mathcal{F}$ can be uncountable, we apply a discretization argument using the covering number. By the definition of the supremum, for any $\epsilon^{\prime}\in\mathbb{R}^{+}$, there exists $f^{*}\in\mathcal{F}$ such that
$\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,Z_{i},Z^{\prime}_{i})=\mathbb{E}_{\sigma}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,Z_{i},Z^{\prime}_{i})+\epsilon^{\prime}.$ (100)
Here we present a stronger result compared to the theorem presented in the main paper, based on the $L_{2}(S_{n})$ pseudometric, which is defined as:
$\|f-f^{\prime}\|_{L_{2}(S_{n})}\coloneqq\sqrt{\frac{1}{n}\sum_{i=1}^{n}|f(X_{i})-f^{\prime}(X_{i})|^{2}}.$
By the definition of the $\epsilon$-covering, for a given $f^{*}\in\mathcal{F}$, there exists $\tilde{f}\in\mathcal{N}(\epsilon,\mathcal{F},L_{2}(S_{2n}))$ such that
$\|f^{*}-\tilde{f}\|_{L_{2}(S_{2n})}\leq\epsilon.$
We first derive the generalization bound using the $L_{2}(S_{n})$ pseudometric, and subsequently upper bound it in terms of the $\|\cdot\|_{\infty}$ norm. By definition, we have
$\mathbb{E}_{\sigma}\max_{\omega}\min_{\omega^{\prime}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f^{*},Z_{i},Z^{\prime}_{i})\leq\mathbb{E}_{\sigma}\max_{\omega}\min_{\omega^{\prime}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(\tilde{f},Z_{i},Z^{\prime}_{i})+2\epsilon.$ (101)
To prove this, we fix $\sigma$ and expand the difference:
$\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\left[(y_{i}-f^{*}(x_{i}))\left(\frac{1+\sigma_{i}}{2}\omega_{i}+\frac{1-\sigma_{i}}{2}\omega^{\prime}_{i}\right)-(y^{\prime}_{i}-f^{*}(x^{\prime}_{i}))\left(\frac{1+\sigma_{i}}{2}\omega^{\prime}_{i}+\frac{1-\sigma_{i}}{2}\omega_{i}\right)\right]$ (102)
$=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\left[(y_{i}-\tilde{f}(x_{i}))\left(\frac{1+\sigma_{i}}{2}\omega_{i}+\frac{1-\sigma_{i}}{2}\omega^{\prime}_{i}\right)-(y^{\prime}_{i}-\tilde{f}(x^{\prime}_{i}))\left(\frac{1+\sigma_{i}}{2}\omega^{\prime}_{i}+\frac{1-\sigma_{i}}{2}\omega_{i}\right)\right]$ (103)
$\quad+\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\left[(\tilde{f}(x_{i})-f^{*}(x_{i}))\left(\frac{1+\sigma_{i}}{2}\omega_{i}+\frac{1-\sigma_{i}}{2}\omega^{\prime}_{i}\right)-(\tilde{f}(x^{\prime}_{i})-f^{*}(x^{\prime}_{i}))\left(\frac{1+\sigma_{i}}{2}\omega^{\prime}_{i}+\frac{1-\sigma_{i}}{2}\omega_{i}\right)\right]$ (104)
$\leq\text{(first term)}+\sqrt{\frac{1}{n}\sum_{i=1}^{n}|\tilde{f}(x_{i})-f^{*}(x_{i})|^{2}}+\sqrt{\frac{1}{n}\sum_{i=1}^{n}|\tilde{f}(x^{\prime}_{i})-f^{*}(x^{\prime}_{i})|^{2}}$ (105)
$\leq\text{(first term)}+2\epsilon,$ (106)
where we used the Cauchy–Schwarz inequality and the bounds $|\omega_{i}|,|\omega^{\prime}_{i}|\leq 1$. Taking expectations over $\sigma$ and maximizing/minimizing over $\omega,\omega^{\prime}$, the result follows.
Thus, we obtain:
$R(\mathcal{F},S_{n},S^{\prime}_{n})\leq\sqrt{\frac{2\log N(\epsilon,\mathcal{F},L_{2}(S_{2n}))}{n}}+2\epsilon+\epsilon^{\prime}.$ (107)
Letting $\epsilon^{\prime}\to 0$, we have
$R(\mathcal{F},S_{n},S^{\prime}_{n})\leq 2\epsilon+\sqrt{\frac{2\log N(\epsilon,\mathcal{F},L_{2}(S_{n}))}{n}}.$ (108)
We can use Eq. equation 108 for the uniform convergence bound, however, to get the refined dependency, we use the chaining technique; Let $\epsilon_{\ell}=2^{-\ell}$ for $\ell=0,1,2,\dots$. Let $\mathcal{F}_{\ell}$ be an $\epsilon_{\ell}$-cover of $\mathcal{F}$ with metric $L_{2}(S_{2n})$, and define $N_{\ell}=|\mathcal{F}_{\ell}|=N(\epsilon_{\ell},\mathcal{F},L_{2}(S_{2n}))$. We may set $\mathcal{F}_{0}=\{0\}$ at scale $\epsilon_{0}=1$.
For each $f\in\mathcal{F}$, we consider $f_{\ell}(f)\in\mathcal{F}_{\ell}$ so that
$\|f-f_{\ell}(f)\|_{L_{2}(S_{2n})}\leq\epsilon_{\ell}.$
Based on the standard chaining technique, which uses $f\in\mathcal{F}$, we consider the following multi-scale decomposition:
$f=(f-f_{L}(f))+\sum_{\ell=1}^{L}(f_{\ell}(f)-g_{\ell-1}(f)).$
By the triangle inequality, we have
$\|f_{\ell}(f)-f_{\ell-1}(f)\|_{L_{2}(S_{n})}\leq\|f_{\ell}(f)-f\|_{L_{2}(S_{n})}+\|f_{\ell-1}(f)-f\|_{L_{2}(S_{n})}\leq 3\epsilon_{\ell}.$
Note that the number of distinct $f_{\ell}(f)-f_{\ell-1}(f)$ is at most $N_{\ell}N_{\ell-1}$.
Similar to the derivation of Eqs. equation 107 and equation 108, regarding $\phi(f,Z_{i},Z^{\prime}_{i})$ as $f(Z_{i})$, we need to care that $2\epsilon$ appears. Therefore, given $\epsilon$-cover for $\mathcal{F}$,
$\|f-f_{\ell}(f)\|_{L_{2}(S_{2n})}\leq 2\epsilon_{\ell}.$
and
$\|f_{\ell}(f)-f_{\ell-1}(f)\|_{L_{2}(S_{n})}\leq\|f_{\ell}(f)-f\|_{L_{2}(S_{n})}+\|f_{\ell-1}(f)-f\|_{L_{2}(S_{n})}\leq(2+4)\epsilon_{\ell}.$
holds if regarding $\phi(f,Z_{i},Z^{\prime}_{i})$ as $f(Z_{i})$ in the above.
Then we can consider the following decomposition;
$R(\mathcal{F},S_{n},S^{\prime}_{n})$ $=\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\left[(f-f_{L}(f))(Z_{i})+\sum_{\ell=1}^{L}(f_{\ell}(f)-f_{\ell-1}(f))(Z_{i})\right]$ (109)
$\leq\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}(f-f_{L}(f))(Z_{i})+\sum_{\ell=1}^{L}\mathbb{E}_{\sigma}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}(f_{\ell}(f)-f_{\ell-1}(f))(Z_{i})$ (110)
${\leq}2\epsilon_{L}+\sum_{\ell=1}^{L}\sup_{f\in\mathcal{F}}\|f_{\ell}(f)-f_{\ell-1}(f)\|_{L_{2}(S_{n})}\sqrt{\frac{2\ln(N_{\ell}N_{\ell-1})}{n}}$ (111)
${\leq}2\epsilon_{L}+6\sum_{\ell=1}^{L}\epsilon_{\ell}\sqrt{\frac{2\ln(N_{\ell}N_{\ell-1})}{n}}$ (112)
$\leq 2\epsilon_{L}+24\sum_{\ell=1}^{L}(\epsilon_{\ell}-\epsilon_{\ell+1})\sqrt{\frac{\ln N_{\ell}}{n}}$ (113)
$\leq 2\epsilon_{L}+24\int_{\epsilon_{L+1}}^{\epsilon_{0}}\frac{\sqrt{\ln N(\epsilon^{\prime},\mathcal{F},L_{2}(S_{n}))}}{\sqrt{n}}d\epsilon^{\prime}.$ (114)
Then for any $\epsilon>0$, we pick $L=\sup\{j:\epsilon_{j}>2\epsilon\}$. This simple $\epsilon_{L+1}\leq 2\epsilon$, thus $\epsilon_{L}\leq 4\epsilon$ holds. By definition of $L$, $\epsilon_{L}>2\epsilon$ and this implies $\epsilon_{L+1}>\epsilon$
Therefore, we have
$R(\mathcal{F},S_{n},S^{\prime}_{n})\leq\inf_{\epsilon\geq 0}\left[8\epsilon+24\int_{\epsilon}^{1}\frac{\sqrt{\ln N(\epsilon^{\prime},\mathcal{F},L_{2}(S_{n}))}}{\sqrt{n}}d\epsilon^{\prime}\right].$ (115)
By definition, we only need to take $\epsilon\leq 1$.
Finally using the fact that $\ln N(\epsilon^{\prime},\mathcal{F},L_{2}(S_{n}))\leq\ln N(\epsilon^{\prime},\mathcal{F},\|\cdot\|_{\infty})$ [wainwright_2019], we obtain the result.
Finally, we remark on the case of $\mathrm{smCE}^{\sigma}$. The dual smooth CE under the dataset $S$ is equivalent to the optimal value of the following optimization problem:
$\mathrm{smCE}^{\sigma}(g,S)=\max_{\{\omega_{i}\}}\frac{1}{n}\sum_{i=1}^{n}(y_{i}-v_{i})\omega_{i}$ (116)
subject to the constraints:
$1\leq\omega_{i}\leq 1,\quad\forall i,\quad|\omega_{i}-\omega_{j}|\leq\frac{1}{4}|g(X_{i})-g(X_{j})|,\quad\forall i,j,$
where $v_{i}=\sigma(g(X_{i}))$, and $\sigma$ denotes the sigmoid function. These constraints can also be rewritten as:
$1\leq\omega_{i}\leq 1,\quad\forall i,\quad|\omega_{i}-\omega_{j}|\leq\frac{1}{4}|\sigma^{-1}(v_{i})-\sigma^{-1}(v_{j})|,\quad\forall i,j.$
By definition, the only difference from the standard smooth CE formulation lies in the constraint of the linear program. Since the resulting problem remains a convex optimization, the same proof techniques developed above can be applied to the dual formulation as well. To carry out a similar analysis, it is necessary to bound the range of the logit function $g$.
C.2 Proof of Lemma 1
Proof.
Since $\phi(f,S)=\frac{1}{n}\sum_{i=1}^{n}\phi(f,z_{i})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f(x_{i}))\omega^{*}_{i}$, where $S=\{z_{i}\}_{i=1}^{n}$ consists of i.i.d. samples and $|(y_{i}-f(x_{i}))\omega^{*}_{i}|\leq 1$, one may wish to apply Hoeffding’s inequality. However, the coefficients $\omega^{*}_{i}$ are obtained as the solution to a linear program that depends on the dataset $S_{\mathrm{te}}$. Consequently, the terms $(y_{i}-f(x_{i}))\omega^{*}_{i}$ are not independent, and Hoeffding’s inequality cannot be applied.
Instead, we employ McDiarmid’s inequality, which requires only the bounded difference property of $\phi(f,S_{\mathrm{te}})$. For completeness, we state McDiarmid’s inequality below:
Lemma 3.
[Boucheron2013] We say that a function $f:\mathcal{X}\to\mathbb{R}$ has the bounded difference property if for some nonnegative constants $c_{1},\dots,c_{n}$,
$\sup_{x_{1},\dots,x_{n},x_{i}^{\prime}\in\mathcal{X}}|f(x_{1},\dots,x_{n})-f(x_{1},\dots,x_{i-1},x^{\prime}_{i},x_{i+1},\dots,x_{n})|\leq c_{i},\quad 1\leq i\leq n.$ (117)
If $X_{1},\dots,X_{n}$ are independent random variables taking values in $\mathcal{X}$ and $f$ has the bounded difference property with constants $c_{1},\dots,c_{n}$, then for any $t\in\mathbb{R}$, we have
$\mathbb{E}\left[e^{t(f(X_{1},\dots,X_{n})-\mathbb{E}[f(X_{1},\dots,X_{n})])}\right]\leq e^{\frac{t^{2}}{8}\sum_{i=1}^{n}c_{i}^{2}}.$ (118)
Moreover
$\mathrm{Pr}\left(f(X_{1},\dots,X_{n})-\mathbb{E}[f(X_{1},\dots,X_{n})]\geq\epsilon\right)\leq e^{\frac{-2\epsilon^{2}}{\sum_{i=1}^{n}c_{i}^{2}}}$ (119)
$\mathrm{Pr}\left(f(X_{1},\dots,X_{n})-\mathbb{E}[f(X_{1},\dots,X_{n})]\leq-\epsilon\right)\leq e^{\frac{-2\epsilon^{2}}{\sum_{i=1}^{n}c_{i}^{2}}}$ (120)
Therefore, we are required to estimate the constants $c_{i}$ in Eq. equation 117. To this end, consider replacing the $i$-th data point $z_{i}$ with $z^{\prime}_{i}$, and let the resulting dataset be denoted by $S^{\prime}_{n}=(z_{1},\dots,z_{i-1},z^{\prime}_{i},z_{i+1},\dots,z_{n})$. We define the following notation:
$\mathrm{smCE}(f,S_{n})=\frac{1}{n}\sum_{j=1}^{n}(y_{j}-f(x_{j}))\omega^{*}_{j},$ (121)
where $\omega^{*}_{j}$ is the solution of Eq. equation 62 given $f$ and $S_{n}$. Similarly, define
$\mathrm{smCE}(f,S^{\prime}_{n})=\frac{1}{n}\sum_{j\neq i}^{n}(y_{j}-f(x_{j}))\omega^{{}^{\prime}*}_{j}+\frac{1}{n}(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{{}^{\prime}*}_{i},$ (122)
where $\omega^{{}^{\prime}*}_{j}$ is the solution of Eq. equation 62 given $f$ and $S^{\prime}_{n}$.
We now evaluate the change in the smooth CE under this replacement:
$\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,S^{\prime}_{n})$ $=\frac{1}{n}\sum_{j=1}^{n}(y_{j}-f(x_{j}))\omega^{*}_{j}-\left[\frac{1}{n}\sum_{j\neq i}^{n}(y_{j}-f(x_{j}))\omega^{{}^{\prime}*}_{j}+\frac{1}{n}(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{{}^{\prime}*}_{i}\right]$ (123)
$\leq\frac{1}{n}\sum_{j=1}^{n}(y_{j}-f(x_{j}))\omega^{*}_{j}-\left[\frac{1}{n}\sum_{j\neq i}^{n}(y_{j}-f(x_{j}))\omega^{*}_{j}+\frac{1}{n}(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{*}_{i}\right]$ (124)
$=\frac{1}{n}(y_{i}-f(x_{i}))\omega^{*}_{i}-\frac{1}{n}(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{*}_{i}$ (125)
$=\frac{1}{n}(y_{i}-y^{\prime}_{i})\omega^{*}_{i}+\frac{1}{n}(f(x^{\prime}_{i})-f(x_{i}))\omega^{*}_{i}\leq\frac{2}{n},$ (126)
where the inequality uses the definition of $\omega^{*}_{j}$ and the bound $|\omega^{*}_{i}|\leq 1$. Similarly, we have:
$\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,S^{\prime}_{n})$ $=\frac{1}{n}\sum_{j=1}^{n}(y_{j}-f(x_{j}))\omega^{*}_{j}-\left[\frac{1}{n}\sum_{j\neq i}^{n}(y_{j}-f(x_{j}))\omega^{{}^{\prime}*}_{j}+\frac{1}{n}(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{{}^{\prime}*}_{i}\right]$ (127)
$\geq\frac{1}{n}\sum_{j=1}^{n}(y_{j}-f(x_{j}))\omega^{{}^{\prime}*}_{j}-\left[\frac{1}{n}\sum_{j\neq i}^{n}(y_{j}-f(x_{j}))\omega^{{}^{\prime}*}_{j}+\frac{1}{n}(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{{}^{\prime}*}_{i}\right]$ (128)
$=\frac{1}{n}(y_{i}-f(x_{i}))\omega^{{}^{\prime}*}_{i}-\frac{1}{n}(y^{\prime}_{i}-f(x^{\prime}_{i}))\omega^{{}^{\prime}*}_{i}$ (129)
$=\frac{1}{n}(y^{\prime}_{i}-y_{i})\omega^{*}_{i}+\frac{1}{n}(f(x^{\prime}_{i})-f(x_{i}))\omega^{{}^{\prime}*}_{i}\geq-\frac{2}{n}.$ (130)
Combining the two results, we obtain:
$\left|\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,S^{\prime}_{n})\right|\leq\frac{2}{n}.$ (131)
Thus, the bounded difference constant $c_{i}$ in McDiarmid’s inequality satisfies:
$\sup_{\{z_{j}\}_{j=1}^{n},\tilde{z}^{\prime}_{i}}\left|\mathrm{smCE}(f,\mathcal{D},S_{n})-\mathrm{smCE}(f,\mathcal{D},S^{\prime}_{n})\right|\leq\frac{1}{n}.$ (132)
Our goal is now to study the uniform stability of the quantity $\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|$. By definition, for any $\epsilon\in\mathbb{R}^{+}$, there exists $g\in\mathcal{F}$ such that
$\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|\leq|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|+\epsilon.$ (133)
Therefore, it suffices to study the stability coefficient $c_{i}$ for $|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|$. Consider the combined dataset $S=S_{\mathrm{te}}\cup S_{\mathrm{tr}}\sim\mathcal{D}^{2n}$, and analyze the effect of replacing a single data point in $S$, which consists of $2n$ i.i.d. samples.
We first consider the case where the replaced data point is from the test set $S_{\mathrm{te}}=\{\tilde{z}_{i}\}_{i=1}^{n}$. Let the perturbed dataset be
$S_{\mathrm{te}}^{\prime}=(\tilde{z}_{1},\dots,\tilde{z}_{i-1},z^{\prime}_{i},\tilde{z}_{i+1},\dots,\tilde{z}_{n}).$
Then, using the triangle inequality and Eq. equation 132, we have:
$|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})|$ $\leq|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{te}}^{\prime})|\leq\frac{2}{n},$ (134)
$|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})|$ $\geq-|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{te}}^{\prime})|\geq-\frac{2}{n}.$ (135)
Combining these, we obtain:
$\left||\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})|\right|\leq\frac{2}{n}.$ (136)
A similar analysis applies when the replaced data point is from the training set $S_{\mathrm{tr}}=\{z_{i}\}_{i=1}^{n}$. Let the perturbed test set be
$S_{\mathrm{tr}}^{\prime}=(z_{1},\dots,z_{i-1},z^{\prime}_{i},z_{i+1},\dots,z_{n}).$
Then,
$\left||\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})|-|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}}^{\prime})|\right|\leq\frac{2}{n}.$ (137)
Therefore, the bounded difference coefficient for each data point in the combined dataset is
$c_{i}=\sup_{\{z_{j}\}_{j=1}^{2n},z^{\prime}_{i}}\left|\phi(g,S_{\mathrm{te}})-\phi(g,S_{\mathrm{tr}})-\left(\phi(g,S_{\mathrm{te}}^{\prime})-\phi(g,S_{\mathrm{tr}})\right)\right|\leq\frac{2}{n},$ (138)
for all $i=1,\dots,2n$.
Applying McDiarmid’s inequality (Eq. equation 119) with these coefficients yields:
$\Pr\left(\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|-\mathbb{E}_{S_{\mathrm{te}},S_{\mathrm{tr}}\sim\mathcal{D}^{2n}}\sup_{f\in\mathcal{F}}|\phi(f,S_{\mathrm{te}})-\phi(f,S_{\mathrm{tr}})|\geq\epsilon\right)\leq\exp(-n\epsilon^{2}/4).$ (139)
C.3 Proof of Lemma 2
Similar to the proof of Lemma 1 in Appendix C.2, we apply McDiarmid’s inequality to evaluate the exponential moment. To this end, we first compute the bounded difference constants in Eq. equation 117. Recall the definition:
$\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,Z_{i},Z^{\prime}_{i}),$ (140)
where
$\phi(f,Z_{i},Z^{\prime}_{i})\coloneqq\left[(y_{i}-f(x_{i}))\left(\tfrac{1+\sigma_{i}}{2}\omega_{i}+\tfrac{1-\sigma_{i}}{2}\omega^{\prime}_{i}\right)-(y^{\prime}_{i}-f(x^{\prime}_{i}))\left(\tfrac{1+\sigma_{i}}{2}\omega^{\prime}_{i}+\tfrac{1-\sigma_{i}}{2}\omega_{i}\right)\right].$ (141)
For clarity, we first analyze the special case where $\sigma_{i}=1$ for all $i\in[n]$; the general case is handled later. Under this assumption, the expression reduces to the difference of smooth CEs:
$\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\phi(f,Z_{i},Z^{\prime}_{i})=\lambda\big(\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,S^{\prime}_{n})\big).$ (142)
Now fix the dataset and flip a single coordinate $\sigma_{j}$ from $1$ to 1-1. Only the $j$-th term is affected. For $\sigma_{j}=1$, the term is
$(y_{j}-f(x_{j}))\omega_{j}-(y^{\prime}_{j}-f(x^{\prime}_{j}))\omega^{\prime}_{j},$
while for $\sigma_{j}=-1$, it becomes
$(y^{\prime}_{j}-f(x^{\prime}_{j}))\omega_{j}-(y_{j}-f(x_{j}))\omega^{\prime}_{j}.$
This corresponds to swapping the training and test inputs in the smooth CE optimization.
Formally, define the datasets:
$\tilde{S}_{n}=((x_{1},y_{1}),\dots,(x^{\prime}_{j},y^{\prime}_{j}),\dots,(x_{n},y_{n})),\quad\tilde{S}^{\prime}_{n}=((x^{\prime}_{1},y^{\prime}_{1}),\dots,(x_{j},y_{j}),\dots,(x^{\prime}_{n},y^{\prime}_{n})).$
Then the exponent, with $\sigma_{i}=1$ for $i\neq j$ and $\sigma_{j}=-1$, can be written as
$\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i\neq j}\phi(f,Z_{i},Z^{\prime}_{i})-\frac{\lambda}{n}\phi(f,Z_{j},Z^{\prime}_{j})=\lambda\big(\mathrm{smCE}(f,\tilde{S}_{n})-\mathrm{smCE}(f,\tilde{S}^{\prime}_{n})\big).$ (143)
Using Eq. equation 132, which bounds the stability of $\mathrm{smCE}$, and combining with Eq. equation 142, we obtain:
$\left|\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\phi(f,Z_{i},Z^{\prime}_{i})-\left(\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i\neq j}\phi(f,Z_{i},Z^{\prime}_{i})-\frac{\lambda}{n}\phi(f,Z_{j},Z^{\prime}_{j})\right)\right|$ (144)
$=\lambda\big|\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,\tilde{S}_{n})\big|+\lambda\big|\mathrm{smCE}(f,S^{\prime}_{n})-\mathrm{smCE}(f,\tilde{S}^{\prime}_{n})\big|$ (145)
$\leq\tfrac{2\lambda}{n}.$ (146)
Since this bound holds for arbitrary $j\in[n]$, we have:
$\sup_{j\in[n]}\!\Big|\!\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\phi(f,Z_{i},Z^{\prime}_{i})\!-\!\Big(\!\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i\neq j}\!\phi(f,Z_{i},Z^{\prime}_{i})\!-\!\frac{\lambda}{n}\phi(f,Z_{j},Z^{\prime}_{j})\!\Big)\!\Big|\!\leq\!\frac{2\lambda}{n}.$ (147)
The same reasoning applies to any fixed realization $\sigma=\sigma|_{\pm}$ of the Rademacher variables (note that $\sigma|_{\pm}$ is a fixed realization and not a random variable). For each such realization, we construct the datasets $S_{\sigma|_{\pm}}$ and $S^{\prime}_{\sigma|_{\pm}}$ used in the smooth CE optimization. Then,
$\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\sigma|_{\pm}\phi(f,Z_{i},Z^{\prime}_{i})=\lambda\left(\mathrm{smCE}(f,S_{\sigma|_{\pm}})-\mathrm{smCE}(f,S^{\prime}_{\sigma|_{\pm}})\right),$ (148)
and changing any single component of $\sigma$ alters the value by at most $\frac{2\lambda}{n}$, as in Eq. equation 147.
Therefore, by the argument in Appendix C.2, the sensitivity coefficients satisfy $c_{i}=\frac{2\lambda}{n}$ for all $i\in[n]$. Substituting this into Lemma 4 in Appendix C.2 yields the desired exponential moment bound.
C.4 Proof of Corollary 1
Proof.
As stated in the main paper, blasiok2023unify proved in Theorem 9.5 that, with probability at least $1-\delta$ over the draw of the test dataset,
$|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{te}})|\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathrm{Lip}_{1}([0,1],[-1,1]))+\sqrt{\frac{\log\frac{2}{\delta}}{2n}}.$ (149)
They further showed that there exists a universal constant $C$ such that
$\mathfrak{R}_{\mathcal{D},n}(\mathrm{Lip}_{1}([0,1],[-1,1]))\leq\frac{C}{\sqrt{n}},$
based on the result of ambroladze2004complexity. See also luxburg2004distance for the derivation of Rademacher complexity of lipshictz functions.
Combining this with Theorem 1 and Corollary 6, and applying the triangle inequality, we obtain:
$\sup_{f}|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ (150)
$\leq\sup_{f}|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{te}})|+\sup_{f}|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|.$ (151)
We now allocate the total failure probability $\delta$ across the two terms using the union bound. Specifically, we set $\delta\to\frac{2}{3}\delta$ in Eq. equation 149, and $\delta\to\frac{1}{3}\delta$ in Theorem 1 or Corollary 6. Then, the sum of the confidence terms becomes:
$\sqrt{\frac{\log\frac{2}{\frac{2}{3}\delta}}{2n}}+2\sqrt{\frac{\log\frac{1}{\frac{1}{3}\delta}}{n}}=\left(2+\frac{1}{\sqrt{2}}\right)\sqrt{\frac{\log\frac{3}{\delta}}{n}}\leq 3\sqrt{\frac{\log\frac{3}{\delta}}{n}}.$ (152)
Thus, we obtain the desired high-probability bound. ∎
Appendix D Analysis based on Rademacher complexity
D.1 Proof of Theorem 2
To simplify the notation, we define
$\phi(f,h,x,y)=h(f(x))\cdot(y-f(x)).$ (153)
Then
$\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})$ (154)
$\leq\sup_{f\in\mathcal{F}}\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})$ (155)
$\leq\sup_{f\in\mathcal{F}}\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\mathbb{E}\phi(f,h,X,Y)-\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})$ (156)
$\leq\sup_{f\in\mathcal{F}}\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\left\{\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})\right\}$ (157)
$\leq\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\left\{\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})\right\}.$ (158)
In the last line, we simply used the swapping of the supremum, which can be shown using the definition of the supremum. (In general, for any $f\in\mathcal{F}$, $\phi(f,h)\leq\sup_{f}^{\prime}\phi(f^{\prime},h)$, which leads to $\sup_{h}\phi(f,h)\leq\sup_{h}\sup_{f}^{\prime}\phi(f^{\prime},h)$. This further leads to $\sup_{f}\sup_{h}\phi(f,h)\leq\sup_{h}\sup_{f}^{\prime}\phi(f^{\prime},h)$.)
Then, we use Mcdiramid’s inequality. The bounded difference for the training dataset is $2/n$, we have the following result; for any $\delta>0$ with probability $1-\delta/2$, we have
$\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})$ (159)
$\leq\mathbb{E}_{S_{\mathrm{tr}}\sim D^{n}}\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (160)
Then by considering the standard symmetrization property, we have the following upper bound
$\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})$ (161)
$\leq 2\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (162)
Finally, by considering the union bound, for any $\delta>0$ with probability $1-\delta$, we have
$\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\Big|\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})\Big|$ (163)
$\leq 2\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (164)
Next, we evaluate the Rademacher complexity term.
For the notational convenience, we define
$T$ $:=\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right],$ (165)
$\eta_{0}$ $:=\{\eta:[0,1]\!\to\!\mathbb{R}\ \mid\ -1\leq\mathrm{Lip}(\eta)\leq 1\}.$
Our goal is to show
$T\leq 2{\mathfrak{R}}_{\mathcal{D},n}(\mathcal{F})+\frac{12}{\sqrt{n}}\int_{0}^{1}\sqrt{\log N\!\bigl(u,\eta_{0},\|\cdot\|_{\infty}\bigr)}du.$ (166)
First, we will consider an $\epsilon$-net of the Lipschitz class, which is independent of data and $\sigma$. Fix geometric scales $\varepsilon_{k}:=2^{-k}$ for $k=0,1,2,\dots$. For each $k$, choose a fixed (data- and $\sigma$-independent) $\varepsilon_{k}$-net $\mathcal{N}_{k}\subset\eta_{0}$ under the uniform norm $\|\cdot\|_{\infty}$, with $|\mathcal{N}_{k}|=N(\varepsilon_{k},\eta_{0},\|\cdot\|_{\infty})$. For every $\eta\in\eta_{0}$ pick a nearest projection $\pi_{k}(\eta)\in\mathcal{N}_{k}$ so that $\|\eta-\pi_{k}(\eta)\|_{\infty}\leq\varepsilon_{k}$. Then, by the triangle inequality,
$\bigl\|\pi_{k}(\eta)-\pi_{k-1}(\eta)\bigr\|_{\infty}\leq\varepsilon_{k}+\varepsilon_{k-1}\leq 2\varepsilon_{k-1}.$ (167)
Consequently,
$\eta=\pi_{0}(\eta)+\sum_{k\geq 1}\bigl(\pi_{k}(\eta)-\pi_{k-1}(\eta)\bigr).$ (168)
We set $A(h)=\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})$. The discretization error by the $k$-th covering is given as
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathrm{Lip}}A(h)-\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}A(h)$ (169)
$=\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathrm{Lip}}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})-\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})$ (170)
$\leq\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathrm{Lip}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f^{\prime},h,X_{i},Y_{i})-\sup_{h\in\mathcal{N}_{k}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f^{\prime},h,X_{i},Y_{i})+2\epsilon^{\prime}$ (171)
$\leq\epsilon_{k}+2\epsilon^{\prime},$ (172)
where $f^{\prime}$ and $\epsilon^{\prime}$ is coming from the definition of the supremum, that is, there exists some $f^{\prime}\in\mathcal{F}$ such that $\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\leq\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f^{\prime},h,X_{i},Y_{i})+\epsilon^{\prime}$ for any $\epsilon^{\prime}\in\real_{+}$. Since we can take $\epsilon^{\prime}$ arbitrarily small, the discretization error is $\epsilon_{k}$.
Next we show that under the $k$-th covering, we can disentangle the complexity of the lipshicit function class by the covering number as follows:
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}A(h)-\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$ (173)
$=\frac{1}{t}\log\exp{t(\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}A(h)-\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}$ (174)
$\leq\frac{1}{t}\log\mathbb{E}_{S}\mathbb{E}_{\sigma}\exp{t(\sup_{h\in\mathcal{N}_{k}}A(h)-\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}$ (175)
$\leq\frac{1}{t}\log\mathbb{E}_{S}\mathbb{E}_{\sigma}\exp{t(\sup_{h\in\mathcal{N}_{k}}A(h)-\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}$ (176)
$\leq\frac{1}{t}\log\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}\exp{t(A(h)-\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}$ (177)
$\leq\frac{1}{t}\log\mathbb{E}_{S}\mathbb{E}_{\sigma}\sum_{h\in\mathcal{N}_{k}}\exp{t(A(h)-\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}.$ (178)
To evaluate this exponential moment, we use the Mcdiramid’s inequality. The bounded difference coefficient is
$\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})-\left(\sup_{f^{\prime}\in\mathcal{F}}\frac{1}{n}\sum_{i\neq j}^{n}\sigma_{i}\phi(f^{\prime},h,X_{i},Y_{i})+\frac{1}{n}\sigma^{\prime}_{j}\phi(f^{\prime},h,X_{j},Y_{j})\right)$ (179)
$\leq\sup_{f\in\mathcal{F}}\frac{1}{n}(\sigma_{j}\phi(f^{\prime},h,X_{j},Y_{j})-\sigma^{\prime}_{j}\phi(f^{\prime},h,X_{j},Y_{j}))\leq\frac{2}{n}.$ (180)
Thus
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}A(h)-\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$ $\leq\frac{1}{t}\log\sum_{h\in\mathcal{N}_{k}}\exp{t^{2}n(\frac{2}{n})/8}$ (181)
$=\frac{1}{t}\log|\mathcal{N}_{k}|+\frac{t}{2n},$ (182)
Next, we analyze $\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$. $\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$ is the Rademacher complexity of $\phi(f,h,X,Y)$. Moreover from Lemma 7.4 in blasiok2023unify, $\phi(f,h,X,Y)$ is 2-Lipshitz with respect to $h$. Thus by Talagrand’s contraction lemma, $\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})$.
Thus, in conclusion, we have
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{N}_{k}}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+\frac{1}{t}\log|\mathcal{N}_{k}|+\frac{t}{2n}.$ (183)
Thus, combining the discretization result
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+\sqrt{\frac{2\log|\mathcal{N}_{k}|}{n}}+\epsilon_{k}.$ (184)
Then, following Dudley’s integral approach as shown in Eq. equation 108, we have
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$ (185)
$\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+\inf_{\epsilon\geq 0}\left[4\epsilon+12\int_{\epsilon}^{1}\frac{\sqrt{\ln N(\varepsilon^{\prime},\eta_{0},\|\cdot\|_{\infty})}}{\sqrt{n}}d\epsilon^{\prime}\right]$ (186)
$\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+\frac{C}{\sqrt{n}},$ (187)
where we used the covering number estimate of the 1-dimensional Lipschitz functions, see wainwright_2019 for the details.
D.2 Sub-optimal result of the Rademacher complexity
The evaluation of $|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ using Rademacher complexity can be carried out in the same manner as in Appendix D.1, except that randomness is now considered over the $2n$ data points from both the test and train sets. Apart from this difference, the argument is identical, and the resulting bound again depends—just as in Appendix D.1—on the complexity of Lipschitz functions through covering arguments, as well as on the complexity of $\mathcal{F}$. This reflects the inherent complexity of the composite function $h(f(X))$ appearing in the definition of smooth CE. Consequently, compared with Theorem 1, where the covering-based proof avoids dependence on the complexity of Lipschitz functions, this approach yields a suboptimal result.
On the other hand, as shown below, it is also possible to derive the Rademacher complexity bound from the covering-number result of Theorem 1. In this case, the dependence on the complexity of Lipschitz functions disappears, but at the cost of introducing an additional $(\log 2n)^{2}$ factor into the Rademacher complexity bound.
Corollary 6.
Under the same assumptions as in Theorem 1, there exist universal constants $\{C_{i}\}$s such that with probability at least $1-\delta$ over the draw of $S_{\mathrm{te}}$ and $S_{\mathrm{tr}}$, we have:
$\sup_{f\in\mathcal{F}}|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|\leq C_{1}\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F})(\log 2n)^{2}+\frac{C_{2}}{\sqrt{n}}+2\sqrt{\frac{\log\delta^{-1}}{n}}.$ (188)
Proof.
We begin by defining the empirical Gaussian complexity of $\mathcal{F}$ as:
$\mathfrak{G}(\mathcal{F},S_{n})=\mathbb{E}_{\mathbf{g}}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}g_{i}f(Z_{i}),$ (189)
where $\mathbf{g}=[g_{1},\dots,g_{n}]$ are independent standard normal random variables, i.e., $g_{i}\sim\mathcal{N}(0,1)$ for all $i\in[n]$.
From the modified version of Sudakov’s minoration inequality (Theorem 12.4 in Zhang_2023), we have:
$\sqrt{\ln\mathcal{N}(\epsilon,\mathcal{F},L_{2}(S_{2n}))}\leq\sqrt{\ln M(\epsilon,\mathcal{F},L_{2}(S_{2n}))}\leq\frac{2\sqrt{2n}\mathfrak{G}_{\mathcal{D},2n}(\mathcal{F})}{\epsilon}+1.$ (190)
This justifies the reason why we use the covering number based on the $L_{2}(S_{2n})$ pseudometric in the derivation of Theorem 1.
As a result, we obtain:
$|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ $\leq\inf_{\epsilon>0}\left[8\epsilon+24\int_{\epsilon}^{1}\left(\frac{2\sqrt{2}\mathfrak{G}_{\mathcal{D},2n}(\mathcal{F})}{\epsilon^{\prime}}+\frac{1}{\sqrt{n}}\right)d\epsilon^{\prime}\right]+\sqrt{\frac{\log(1/\delta)}{n}}$ (191)
$\leq\inf_{\epsilon>0}\left[8\epsilon+48\sqrt{2}\mathfrak{G}_{\mathcal{D},2n}(\mathcal{F})\log\left(\frac{1}{\epsilon}\right)+\frac{24}{\sqrt{n}}\right]+\sqrt{\frac{\log(1/\delta)}{n}}$ (192)
$\leq\frac{8}{\sqrt{2n}}+24\sqrt{2}\mathfrak{G}_{\mathcal{D},2n}(\mathcal{F})\log(2n)+\frac{24}{\sqrt{2n}}+\sqrt{\frac{\log(1/\delta)}{n}},$ (193)
where we have set $\epsilon=1/\sqrt{2n}$ in the final inequality.
Furthermore, from Lemma 4 in bartlett2002rademacher, there exist universal constants $c$ and $C$ such that:
$c\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F})\leq\mathfrak{G}_{\mathcal{D},2n}(\mathcal{F})\leq C\log(2n)\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F}).$ (194)
Combining the above results, we conclude:
$|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|\leq 24\sqrt{2}C\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F})(\log 2n)^{2}+\frac{32}{\sqrt{2n}}+2\sqrt{\frac{\log(1/\delta)}{n}}.$ (195)
Appendix E Proofs of the gradient boosting tree
E.1 Quadratic upper bound of the boosting
The iterative minimization problem is given as
$\min_{\psi_{\theta}}\mathcal{L}_{n}\big(g^{(t)}-w_{t}\psi_{\theta}(x)\big),$ (196)
and by considering the quadratic upper bound for this
$\min_{\psi_{\theta}}\Big(-\langle\nabla\mathcal{L}_{n}(g^{(t)}),w_{t}\psi_{\theta}\rangle+\frac{M}{2}\|w_{t}\psi_{\theta}\|^{2}\Big)\Leftrightarrow\min_{\psi_{\theta}}\|Mw_{t}\psi_{\theta}-\nabla\|^{2}.$ (197)
E.2 From weak learnability to functional gradient
We use the empirical inner product and norms
$\langle a,b\rangle_{L_{2}(S_{n})}:=\frac{1}{n}\sum_{i=1}^{n}a_{i}b_{i},\quad\|f\|_{L_{2}(S_{n})}^{2}:=\langle f,f\rangle_{L_{2}(S_{n})},\quad\|v\|_{L_{1}(S_{n})}:=\frac{1}{n}\sum_{i=1}^{n}|v_{i}|.$
Consider the binary cross-entropy with logit $z=g(X)$, i.e.,
$\ell_{\rm ent}(\tilde{y},z)\ =\ -\tilde{y}\log\sigma(z)\ -\ (1-\tilde{y})\log\bigl(1-\sigma(z)\bigr),\quad\sigma(z)=\frac{1}{1+e^{-z}}.$
Its per-sample gradient w.r.t. the logit is
$\nabla_{i}\ :=\ \frac{\partial}{\partial z}\ell_{\rm ent}(\tilde{y}_{i},z)\Big|_{z=g(X_{i})}\ =\ \sigma\bigl(g(X_{i})\bigr)\ -\ \tilde{y}_{i}.$
Let $v=(\nabla_{1},\ldots,\nabla_{n})$.
Lemma 4.
Under the same assumptions in Theorem 3, the following holds:
$\mathrm{smCE}^{\sigma}\big(g^{(t)},S_{n}\big)\leq\|v\|_{L_{1}(S_{n})}\leq\frac{1}{\gamma B}\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}+\frac{Mw_{t}}{2\gamma}B.$ (198)
Proof.
Introduce $y_{i}:=2\tilde{y}_{i}-1\in\{\pm 1\}$. Then
$\nabla_{i}=\sigma\bigl(g(X_{i})\bigr)-\tilde{y}_{i}=-y_{i}\sigma\bigl(-y_{i}g(X_{i})\bigr):=-y_{i}q_{i},\quad q_{i}\in(0,1),$
hence $|\nabla_{i}|=q_{i}$. Let $w_{i}:=\dfrac{|\nabla_{i}|}{\sum_{j=1}^{n}|\nabla_{j}|}\in\Delta_{n}$. By Assumption 2 there exists $\psi^{\star}\in\Psi$ such that $\sum_{i=1}^{n}w_{i}y_{i}\psi^{\star}(X_{i})\geq\gamma B$. Therefore
$\big\langle v,\psi^{\star}\big\rangle_{L_{2}(S_{n})}=\frac{1}{n}\sum_{i=1}^{n}(-y_{i}|\nabla_{i}|)\psi^{\star}(X_{i})=-\frac{1}{n}\Big(\sum_{i=1}^{n}|\nabla_{i}|\Big)\Big(\sum_{i=1}^{n}w_{i}y_{i}\psi^{\star}(X_{i})\Big)\ \leq\ -\gamma B\|v\|_{L_{1}(S_{n})}.$
Since $\Psi$ is sign-flip closed, $\psi^{\star}\in\Psi$, and
$\big\langle v,-\psi^{\star}\big\rangle_{L_{2}(S_{n})}\geq\ \gamma B\|v\|_{L_{1}(S_{n})}.$ (199)
Note that
$\|Mw_{t}\psi_{\theta_{t}}-\nabla\|_{L_{2}(S_{n})}^{2}\leq\|Mw_{t}(-\psi^{\star})-\nabla_{t}\|_{L_{2}(S_{n})}^{2},$ (200)
which implies that
$-2Mw_{t}\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}\leq-2Mw_{t}\big\langle(-\psi^{\star}),\ \nabla_{t}\big\rangle_{L_{2}(S_{n})}+M^{2}w_{t}^{2}\|\psi^{\star}\|^{2}_{L_{2}(S_{n})},$ (201)
and we obtain
$\langle(-\psi^{\star}),\ \nabla_{t}\big\rangle_{L_{2}(S_{n})}\leq\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}+\frac{Mw_{t}}{2}B^{2}.$ (202)
Then combining Eq. equation 199 and equation 202 we obtain
$\gamma B\|v\|_{L_{1}(S_{n})}\leq\big\langle\psi_{t},\nabla_{t}\big\rangle_{L_{2}(S_{n})}+\frac{Mw_{t}}{2}B^{2}.$ (203)
By definition, we have
$\mathrm{smCE}^{\sigma}\big(g^{(t)},S_{n}\big)\leq\|v\|_{L_{1}(S_{n})}.$ (204)
Since $M=1/4$, combining the above two inequalities, we obtain the result. ∎
E.3 Proof of Theorem 3
We consider the Taylor expansion of the objective
$L_{n}(g^{(t+1)})$ $\leq L_{n}(g^{(t)})-w_{t}\langle\nabla,\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}+\frac{M}{2}w_{t}^{2}\|\psi_{\theta_{t}}\|_{L_{2}(S_{n})}^{2}.$ (205)
Here we use Lemma 5 in Appendix E.4. The lemma is the Pythagorean relation of the projection of the function gradient. In the lemma, $w_{t}M\to w$ and $\nabla\to x$, and $p\to\psi_{\theta_{t}}$ corresponding to the GBT step. From Eq. equation 213, setting $q=0$ and , we get
$\langle\psi_{\theta_{t}},w_{t}M\psi_{\theta_{t}}-\nabla_{t}\rangle_{L_{2}(S_{n})}\leq 0,$ (206)
and this leads to
$Mw_{t}\|\psi_{\theta_{t}}\|_{n}^{2}\leq\langle\psi_{\theta_{t}},\nabla_{t}\rangle_{L_{2}(S_{n})}.$ (207)
Finally we have
$L_{n}(g^{(t+1)})$ $\leq L_{n}(g^{(t)})-\frac{w_{t}}{2}\langle\nabla_{t},\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}.$ (208)
By summing up from 0 to T1T-1, we have
$\frac{1}{2}\sum_{t=0}^{T-1}w_{t}\langle\nabla_{t},\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}$ $\leq L_{n}(g^{(0)})-L_{n}(g^{(T)}).$ (209)
If we use the constant stepsize $w_{t}=w$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\langle\nabla_{t},\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}$ $\leq\frac{2\big(L_{n}(g^{(0)})-L_{n}(g^{(T)})\big)}{wT}.$ (210)
Combining Lemma 4, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathrm{smCE}^{\sigma}\big(g^{(t)},S_{n}\big)\leq\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla_{t}\|_{L_{1}(S_{n})}$ $\leq\frac{1}{2\gamma B}\cdot\frac{1}{T}\sum_{t=0}^{T-1}\langle\nabla_{t},\psi_{\theta_{t}}\rangle_{L_{2}(S_{n})}+\frac{wMB}{2\gamma}$ (211)
$\leq\frac{1}{2\gamma B}\cdot\frac{2\big(L_{n}(g^{(0)})-L_{n}(g^{(T)})\big)}{wT}+\frac{wMB}{2\gamma}.$ (212)
Since the cross entropy loss is always positive, we drop $L_{n}(g^{(T)})$ and obtain the result.
E.4 Auxiliary lemmas
Setting.
Let $X\subset\mathbb{R}^{d}$ be a nonempty closed convex set and fix a scalar weight $w>0$. Consider
$f(z)\ :=\ \tfrac{1}{2}\|wz-x\|_{2}^{2}\ =\ \tfrac{w^{2}}{2}\big\|z-\tfrac{x}{w}\big\|_{2}^{2}\ +\ \mathrm{const},\qquad x\in\mathbb{R}^{d},$
and let
[…]
Equivalently, with $y_{0}:=x/w$, $p$ is the Euclidean projection $p=\Pi_{X}(y_{0})$.
Following is the alternative lemma of Lemma 3.1 in bubeck2015convex,
Lemma 5.
For all $q\in X$, the following relation holds.
$\langle p-q,w^{2}p-wx\rangle\leq 0.$
Equivalently, since $w^{2}p-wx=w^{2}(p-y_{0})$, we have
$\langle p-q,p-y_{0}\rangle\leq 0.$ (213)
Proof. The objective $f$ is convex and differentiable with gradient
$\nabla f(z)\ =\ w^{2}z-wx.$
For convex differentiable optimization over a convex set $X$, the necessary and sufficient first-order condition at a minimizer $p$ is
$\langle\nabla f(p),q-p\rangle\geq 0\qquad(\forall q\in X).$
Rewriting gives $\langle\nabla f(p),p-q\rangle\leq 0$, i.e. $\langle p-q,w^{2}p-wx\rangle\leq 0$. $\square$
E.5 Summary of the gradient boosting tree algorithm
We remark that the regions of a binary tree are defined as follows: Given a node associated with some region $\tilde{R}\subset\mathcal{X}\subset\mathbb{R}^{d}$, we split the region using a splitting threshold $s\in\mathbb{R}$ and a splitting dimension $j\in[d]$ into two child nodes, $\tilde{R}_{L}$ and $\tilde{R}_{R}$, which respectively correspond to the regions $\tilde{R}_{L}:=\{x\in\tilde{R}\mid x_{j}\leq s\}$ and $\tilde{R}_{R}:=\{x\in\tilde{R}\mid x_{j}>s\}$. By recursively applying this procedure, we obtain partitions $\{R_{j}\}_{j=1}^{J}$, where each region corresponds to a hyperrectangular region in $\mathbb{R}^{d}$.
The set of all such trees defines the following function class:
$\mathcal{T}=\left\{x\mapsto\sum_{j=1}^{J}c_{j}\cdot\mathbbm{1}_{\{x\in R_{j}\}}\middle|J\leq 2^{m},R_{j}\text{ disjoint},c_{j}\in\mathbb{R}\right\}.$
and $|C_{j}|$ is bounded by $B$
Let $\nabla_{t,i}=\nabla_{g}\ell(\sigma(g^{(t)}(X_{i})),Y_{i})$ denote the functional gradient at the $i$-th training point.
Algorithm 1 Gradient Boosting Tree with Cross-Entropy Loss
0: Training data $S_{\mathrm{tr}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, base learner $\psi_{\theta}\in\mathcal{T}$, loss function $\ell_{\mathrm{ent}}$
0: Final logit function $g^{(T)}(x)$
1: Initialize $g^{(0)}(x)=0$
2:for $t=0,1,\dots,T-1$ do
3:  Compute functional gradients:
$\nabla_{t,i}=\nabla_{g}\ell_{\mathrm{ent}}(\sigma(g^{(t)}(X_{i})),Y_{i}),\quad i=1,\dots,n$
4:  (Optionally set step size $w_{t}$)
5:  Solve: $\theta_{t}=\operatorname*{arg\,min}_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}|Mw_{t}\psi_{\theta}(X_{i})-\nabla_{t,i}|^{2}$
6:  Update the model:
$g^{(t+1)}(x)=g^{(t)}(x)-w_{t}\psi_{\theta_{t}}(x)$
7:end for
8:return $g^{(T)}(x)$ or $\bar{g}^{(T)}(x)$
E.6 Proof of Corollary 2
We first remark that the following relationship holds:
$\mathrm{smCE}(f,S_{\mathrm{tr}})\leq\mathrm{smCE}^{\sigma}(g,S_{\mathrm{tr}}).$ (214)
The proof is nearly identical to that of Lemma 4.7 in blasiok2023when. By replacing the population expectation over $\mathcal{D}$ with the empirical expectation over the sample $S_{\mathrm{tr}}$, we obtain the result.
Therefore, an upper bound on the dual smooth CE over the training dataset directly implies an upper bound on the standard smooth CE. By combining this with our developed generalization bounds, we obtain the desired result.
We now evaluate the Rademacher complexity. We express the function class of $\bar{g}^{(T)}=\frac{1}{n}\sum_{t=0}^{T-1}g^{t}$ as $\mathcal{G}$.
Since $\mathcal{F}=\sigma\circ\mathcal{G}$ and $\sigma$ is $1/4$-Lipschitz, Talagrand’s contraction lemma yields:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})=\mathfrak{R}_{\mathcal{D},n}(\sigma\circ\mathcal{G})\leq\frac{1}{4}\mathfrak{R}_{\mathcal{D},n}(\mathcal{G}).$ (215)
We now derive an upper bound on the Rademacher complexity of $\mathcal{G}$. Using its sub-additive property, we have:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{G})\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathfrak{R}_{\mathcal{D},n}(\{g{(t)}\}).$ (216)
where $\{g{(t)}\}$ is the function class that after $t$-step GBT algorithms.
Then
$\mathfrak{R}_{\mathcal{D},n}(\{g^{(t)}\})\leq wt\mathfrak{R}_{\mathcal{D},n}(\mathcal{T}),$ (217)
where $\mathcal{T}$ denotes the class of regression trees.
Furthermore, by Theorem 6.25 in Zhang_2023, the Rademacher complexity is upper bounded by the covering number:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{T})\leq\inf_{\epsilon>0}\left[4\epsilon+12\int_{\epsilon}^{1}\sqrt{\frac{\ln N(\epsilon^{\prime},\mathcal{T},L_{2}(S_{n}))}{n}}d\epsilon^{\prime}\right].$ (218)
To proceed, we upper bound the covering number for binary regression trees. From Appendix B in klusowski2024large, we have:
$\ln N(\epsilon,\mathcal{T},L_{2}(S_{n}))\leq\ln(nd)^{2^{m}}\left(\frac{C}{\epsilon^{2}}\right)^{2^{m+1}}.$ (219)
Therefore, there exists a universal constant $C>0$ such that:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{T})\leq C\sqrt{\frac{2^{m}\log(nd)}{n}}.$ (220)
In conclusion, we have that
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})\leq CwT\sqrt{\frac{2^{m}\log(nd)}{n}}.$ (221)
Finally, substituting this into Theorem 2, we obtain the result.
$\mathrm{smCE}^{\sigma}\big(\frac{1}{T}\sum_{t=0}^{T-1}g^{(t)},S_{n}\big)\leq$ $\frac{L_{n}(g^{(0)})}{\gamma BwT}+\frac{wMB}{2\gamma}+wTC\sqrt{\frac{2^{m}\log(nd)}{n}}\!+\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (222)
E.7 Misclassification rate
Let $S_{n}$ denote the empirical distribution of the training dataset. From standard margin-based generalization theory, for any $\rho>0$, with probability at least $1-\delta$, the following inequality holds:
$P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq P_{(X,Y)\sim S_{n}}[(2Y-1)\bar{g}^{(T)}(X)\leq\rho]+\frac{2}{\rho}\mathfrak{R}_{\mathcal{D},n}(\mathcal{G})+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}.$ (223)
According to nitanda18a, nitanda2019gradient, the empirical margin error can be upper-bounded by the functional gradient as follows:
$P_{(X,Y)\sim S_{n}}[(2Y-1)\bar{g}^{(T)}(X)\leq\rho]\leq(1+e^{\rho})\|\nabla_{g}\ell_{\mathrm{ent}}(\bar{g}^{(T)}(X),Y)\|_{L_{1}(S_{n})}.$ (224)
Combining the above results, we obtain:
$P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq(1+e^{\rho})\left(\frac{L_{n}(g^{(0)})}{\gamma BwT}+\frac{wB}{8\gamma}\right)+\frac{CwT}{\rho}\sqrt{\frac{2^{J}\log(nd)}{n}}+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}.$ (225)
This result illustrates a trade-off with respect to $wT$, which balances the training misclassification rate and the complexity of the hypothesis class.
Appendix F Proofs for the kernel boosting
F.1 Proof of Eq. (23)
From the margin assumption, the following property holds:
$\phi(x_{i})>\gamma\quad\text{if }y_{i}=1,\quad-\phi(x_{i})>\gamma\quad\text{if }y_{i}=0,\quad\text{for all }(x_{i},y_{i})\in S_{\mathrm{tr}}.$
Since $\phi\in\mathcal{H}$ implies $\phi\in\mathcal{H}$, we define $\tilde{\phi}:=-\phi$. Then, the margin assumption implies:
$\tilde{\phi}(x_{i})<-\gamma\quad\text{if }y_{i}=1,\quad\tilde{\phi}(x_{i})>\gamma\quad\text{if }y_{i}=0,\quad\text{for all }(x_{i},y_{i})\in S_{\mathrm{tr}}.$
We now analyze the gradient in the RKHS under this construction:
$\left\|\mathcal{T}_{k}\nabla_{g}L_{n}(g)\right\|_{\mathcal{H}}$ $=\left\langle\mathcal{T}_{k}\nabla_{g}L_{n}(g),\frac{\mathcal{T}_{k}\nabla_{g}L_{n}(g)}{\|\mathcal{T}_{k}\nabla_{g}L_{n}(g)\|_{\mathcal{H}}}\right\rangle_{\mathcal{H}}$ (226)
$=\sup_{\phi\in\mathcal{H},\|\phi\|_{\mathcal{H}}\leq 1}\left\langle\mathcal{T}_{k}\nabla_{g}L_{n}(g),\phi\right\rangle_{\mathcal{H}}$ (227)
$\geq\left\langle\nabla_{g}L_{n}(g),\tilde{\phi}\right\rangle_{L_{2}(S_{n})}$ (228)
$=\frac{1}{n}\sum_{i=1}^{n}\nabla_{g}\ell(g(X_{i}),Y_{i})\cdot\tilde{\phi}(X_{i})$ (229)
$=\frac{1}{n}\sum_{i=1}^{n}(\sigma(g(X_{i}))-Y_{i})\cdot\tilde{\phi}(X_{i}),$ (230)
where the last equality follows from the functional gradient.
Using the margin assumption, we obtain:
$\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\sigma(g(X_{i})))\cdot(-\tilde{\phi}(X_{i}))\geq\frac{\gamma}{n}\sum_{i=1}^{n}\left|\nabla_{g}\ell(g(X_{i}),Y_{i})\right|.$ (231)
This completes the proof.
F.2 Proof of Theorem 4
Assume that the loss function $L(g)$ is $M$-Lipschitz smooth with respect to $g$, meaning that for all $g,g^{\prime}$,
$L(g^{\prime})\leq L(g)+\langle\nabla L(g),g^{\prime}-g\rangle_{L_{2}(S_{n})}+\frac{M}{2}\|g^{\prime}-g\|_{L_{2}(S_{n})}^{2}.$ (232)
Plugging in $g=g^{(t)}$ and $g^{\prime}=g^{(t+1)}=g^{(t)}-w_{t}\mathcal{T}_{k}\nabla_{g}L(g^{(t)})$, we obtain:
$L(g^{(t+1)})$ $\leq L(g^{(t)})-w_{t}\langle\nabla L(g^{(t)}),\mathcal{T}_{k}\nabla_{g}L(g^{(t)})\rangle_{L_{2}(S_{n})}$ (233)
$\quad+\frac{M}{2}w_{t}^{2}\|\mathcal{T}_{k}\nabla_{g}L(g^{(t)})\|_{L_{2}(S_{n})}^{2}$ (234)
$\leq L(g^{(t)})-w_{t}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}+\frac{M\Lambda}{2}w_{t}^{2}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}$ (235)
$=L(g^{(t)})-w_{t}\left(1-\frac{w_{t}M\Lambda}{2}\right)\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}$ (236)
$\leq L(g^{(t)})-\frac{w_{t}}{2}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2},$ (237)
where the last inequality uses $w_{t}\leq\frac{1}{M\Lambda}$.
Summing over $t=0,\ldots,T-1$, we get:
$\sum_{t=0}^{T-1}\frac{1}{2}w_{t}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}\leq L_{n}(g^{(0)}).$ (238)
In particular, using constant step size $w_{t}=w$, we obtain:
$\frac{1}{T}\sum_{t=0}^{T-1}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}\leq\frac{2}{wT}L_{n}(g^{(0)}).$ (239)
From Eq. equation 226, we have:
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla_{g}\ell(g^{(t)}(X),Y)\right\|_{L_{1}(S_{n})}^{2}\leq\frac{2}{wT}L_{n}(g^{(0)}).$ (240)
Then, using Jensen’s inequality and the definition of $\bar{g}^{(T)}:=\frac{1}{T}\sum_{t=0}^{T-1}g^{(t)}$, we get:
$\left\|\nabla_{g}\ell(\bar{g}^{(T)}(X),Y)\right\|_{L_{1}(S_{n})}^{2}\leq\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla_{g}\ell(g^{(t)}(X),Y)\right\|_{L_{1}(S_{n})}^{2}.$ (241)
Combining these gives:
$\left\|\nabla_{g}\ell(\bar{g}^{(T)}(X),Y)\right\|_{L_{1}(S_{n})}\leq\sqrt{\frac{2}{wT}L_{n}(g^{(0)})}.$ (242)
Generalization Bound via Rademacher Complexity.
We now evaluate the Rademacher complexity. We express the function class of $\bar{g}^{(T)}=\frac{1}{n}\sum_{t=0}^{T-1}g^{t}$ as $\mathcal{G}$. Since $\mathcal{F}=\sigma\circ\mathcal{G}$ and $\sigma$ is $1/4$-Lipschitz, Talagrand’s contraction lemma yields:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})=\mathfrak{R}_{\mathcal{D},n}(\sigma\circ\mathcal{G})\leq\frac{1}{4}\mathfrak{R}_{\mathcal{D},n}(\mathcal{G})\leq\frac{\alpha}{4}\sqrt{\frac{\Lambda}{n}},$ (243)
where $\alpha$ is an upper bound on the RKHS norm of $\mathcal{G}$, which we estimate below. The final inequality follows from the standard Rademacher complexity estimate of the RKHS, see Mohri for the details.
Recall the recursion:
$g^{(T)}=g^{(0)}-w\sum_{t=0}^{T-1}\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)}).$ (244)
Then:
$\|g^{(T)}\|_{\mathcal{H}}$ $\leq\|g^{(0)}\|_{\mathcal{H}}+\left\|w\sum_{t=0}^{T-1}\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\right\|_{\mathcal{H}}.$ (245)
Applying Jensen’s inequality:
$\|w\sum_{t=0}^{T-1}\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}\leq\frac{1}{T}\sum_{t=0}^{T-1}\|wT\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}$ $\leq w^{2}T^{2}\frac{1}{T}\sum_{t=0}^{T-1}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}$ (246)
$\leq w^{2}T^{2}\frac{2}{wT}L_{n}(g^{(0)})$ (247)
$\leq 2wTL_{n}(g^{(0)})$ (248)
Assuming $\|g^{(0)}\|_{\mathcal{H}}\leq\Lambda^{\prime}$, we obtain:
$\|g^{(T)}\|_{\mathcal{H}}\leq\Lambda^{\prime}+\sqrt{2wTL_{n}(g^{(0)})}.$ (249)
Absorbing $\sqrt{2L_{n}(g^{(0)})}$ into a universal constant. When considering the norm of $\bar{g}^{(T)}$, it is also $\mathcal{O}(\Lambda^{\prime}+\sqrt{wT})$. Thus, we conclude $\alpha=\mathcal{O}(\Lambda^{\prime}+\sqrt{wT})$. Substituting this into the Rademacher bound yields the final generalization result.
F.3 Proof of Misclassification rate
Let $S_{n}$ denote the empirical distribution of the training dataset. From standard margin-based generalization theory, for any $\rho>0$, with probability at least $1-\delta$, the following inequality holds:
$P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq P_{(X,Y)\sim S_{n}}[(2Y-1)\bar{g}^{(T)}(X)\leq\rho]+\frac{2}{\rho}\mathfrak{R}_{\mathcal{D},n}(\mathcal{G})+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}.$ (250)
According to nitanda18a, nitanda2019gradient, the empirical margin error can be upper-bounded by the functional gradient as follows:
$P_{(X,Y)\sim S_{n}}[(2Y-1)\bar{g}^{(T)}(X)\leq\rho]\leq(1+e^{\rho})\|\nabla_{g}\ell_{\mathrm{ent}}(\bar{g}^{(T)}(X),Y)\|_{L_{1}(S_{n})}.$ (251)
Combining the above results, we obtain:
$P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq\frac{(1+e^{\rho})}{\gamma}\sqrt{\frac{L_{n}(g^{(0)})}{wT}}+\frac{2(\Lambda^{\prime}+\sqrt{2wTL_{n}(g^{(0)})})}{\rho}\sqrt{\frac{\Lambda}{n}}+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}.$ (252)
This result illustrates a trade-off with respect to $wT$, which balances the training misclassification rate and the complexity of the hypothesis class.
F.4 Proof of Corollary 4
We first remark that both the smooth calibration error and the misclassification rate exhibit similar asymptotic behavior with respect to $n$ and $wT$. By substituting the given hyperparameters into the bound presented in Theorem 4, we obtain the desired result.
Appendix G Formal settings of the two-layer neural network
G.1 Formal results for the smooth CE
Here we provide the formal setting of the results from nitanda2019gradient. We assume $Y=\{\pm 1\}$ and denote by $\nu$ the true probability measure on $\mathcal{X}\times\mathcal{Y}$, and by $\nu_{n}$ the empirical measure based on samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$ drawn independently from $\nu$, i.e.,
$d\nu_{n}(X,Y)=\frac{1}{n}\sum_{i=1}^{n}\delta_{(X_{i},Y_{i})}(X,Y)dXdY,$
where $\delta$ denotes the Dirac delta function. In the main paper, we consider labels in $\tilde{Y}=\{0,1\}$; here, we convert them to $Y=2\tilde{Y}-1$, so the distribution $\mathcal{D}$ over $\mathcal{X}\times\tilde{Y}$ becomes $\mu$ over $\mathcal{X}\times Y$.
The marginal distributions of $\nu$ and $\nu_{n}$ over $\mathcal{X}$ are denoted by $\nu_{\mathcal{X}}$ and $\nu_{\mathcal{X},n}$, respectively. For $s\in\mathbb{R}$ and $y\in\mathcal{Y}$, let $\ell(s,y)$ denote the logistic loss:
$\ell(s,y)=\log(1+\exp(-ys)).$
We remark that given a logit value $s$, the predicted probability of $Y=1$ is $p=\sigma(s)$, and the loss can also be written as $\ell(s,y)=\frac{y+1}{2}\log p+\frac{1-y}{2}\log(1-p).$ Replacing $Y$ with $\tilde{Y}$ recovers the standard cross-entropy loss: $\ell(s,\tilde{y})=\tilde{y}\log p+(1-\tilde{y})\log(1-p).$
The empirical objective to be minimized is defined as
$L_{n}(\theta)\coloneqq\mathbb{E}_{(X,Y)\sim\nu_{n}}\left[\ell(g_{\theta}(X),Y)\right]=\frac{1}{n}\sum_{i=1}^{n}\ell(g_{\theta}(X_{i}),Y_{i}),$
where $g_{\theta}:\mathcal{X}\rightarrow\mathbb{R}$ is a two-layer neural network parameterized by $\theta=(\theta_{r})_{r=1}^{m}$. When treating the function $g_{\theta}$ as the variable of the objective, we also write $L_{n}(g_{\theta})\coloneqq L_{n}(\theta)$.
We now introduce the following formal assumptions.
Assumption 1.
  • (A1)
    Assume that $\mathrm{supp}(\nu_{\mathcal{X}})\subset\{x\in\mathcal{X}\mid\|x\|_{2}\leq 1\}$. Let $\sigma$ be a $C^{2}$-class function, and there exist constants $K_{1},K_{2}>0$ such that
    $\|\sigma^{\prime}\|_{\infty}\leq K_{1}\quad\text{and}\quad\|\sigma^{\prime\prime}\|_{\infty}\leq K_{2}.$
  • (A2)
    A distribution $\mu_{0}$ on $\mathbb{R}^{d}$, used for the initialization of $\theta_{r}$, has a sub-Gaussian tail bound: there exist constants $A,b>0$ such that
    $\mathbb{P}_{\theta^{(0)}\sim\mu_{0}}\left[\|\theta^{(0)}\|_{2}\geq t\right]\leq A\exp(-bt^{2}).$
  • (A3)
    Assume that the number of hidden units $m\in\mathbb{Z}_{+}$ is even. The constant parameters $(a_{r})_{r=1}^{m}$ and parameters $\theta^{(0)}=(\theta_{r}^{(0)})_{r=1}^{m}$ are initialized symmetrically as follows:
    $a_{r}=1\quad\text{for }r\in\left\{1,\dots,\frac{m}{2}\right\},\quad a_{r}=-1\quad\text{for }r\in\left\{\frac{m}{2}+1,\dots,m\right\},$
    and
    $\theta_{r}^{(0)}=\theta_{r+\frac{m}{2}}^{(0)}\quad\text{for }r\in\left\{1,\dots,\frac{m}{2}\right\},$
    where the initial parameters $(\theta_{r}^{(0)})_{r=1}^{m/2}$ are independently drawn from the distribution $\mu_{0}$.
  • (A4)
    Assume that there exist $\gamma>0$ and a measurable function $v:\mathbb{R}^{d}\to\{w\in\mathbb{R}^{d}\mid\|w\|_{2}\leq 1\}$ such that the following inequality holds for all $(x,y)\in\mathrm{supp}(\nu)\subset\mathcal{X}\times\mathcal{Y}$:
    $y\left\langle\partial_{\theta}\sigma(\theta^{(0)}\cdot x),v(\theta^{(0)})\right\rangle_{L^{2}(\mu_{0})}=y\mathbb{E}_{\theta^{(0)}\sim\mu_{0}}\left[\partial_{\theta}\sigma(\theta^{(0)\top}x)\cdot v(\theta^{(0)})\right]\geq\gamma.$
Remark.
Clearly, many activation functions (sigmoid, tanh, and smooth approximations of ReLU such as swish) satisfy assumption (A1). Typical distributions, including the Gaussian distribution, satisfy (A2). The purpose of the symmetrized initialization (A3) is to uniformly bound the initial value of the loss function $L_{n}(\theta^{(0)})$ over the number of hidden units $m$. This initialization leads to $f_{\theta^{(0)}}(x)=0$, resulting in $L_{n}(\theta^{(0)})=\log(2)$. Assumption (A4) implies the separability of a dataset using the neural tangent kernel. We next discuss the validity of this assumption.
Theorem 6 (Global Convergence, Theorem 2 in nitanda2019gradient).
Suppose Assumption 1 holds. We set constant $K$ as
$K=K_{1}^{4}+2K_{1}^{2}K_{2}+K_{1}^{4}K_{2}^{2}.$ (253)
For all $\beta\in[0,1)$, $\delta\in(0,1)$, and $m\in\mathbb{Z}_{+}$ such that
$m\geq\frac{16K_{1}^{2}}{\gamma^{2}}\log\frac{2n}{\delta},$
consider gradient descent (6) with learning rate
$0<w\leq\min\left\{\frac{1}{m^{\beta}},\frac{4m^{2\beta-1}}{K_{1}^{2}+K_{2}}\right\},$
and the number of iterations
$T\leq\left\lfloor\frac{m\gamma^{2}}{32wK_{2}^{2}\log(2)}\right\rfloor.$
Then, with probability at least $1-\delta$ over the random initialization, we have:
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla_{g}L_{n}(g_{\theta^{(t)}})\right\|^{2}_{L^{1}(\nu_{\mathcal{X},n})}\leq\frac{16\log(2)}{\gamma^{2}T}\left(\frac{m^{2\beta-1}}{w}+K\right).$
where we define the $L^{1}$-norm of the functional gradient as
$\left\|\nabla_{g}L_{n}(g_{\theta})\right\|_{L^{1}(\nu_{\mathcal{X},n})}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\left|\partial_{g}\ell(g_{\theta}(X_{i}),Y_{i})\right|=\frac{1}{2n}\sum_{i=1}^{n}\left|Y_{i}-2p_{\theta}(Y=1\mid x_{i})+1\right|.$
The result above corresponds to Eq. equation 27. This bound provides a guarantee on the training smooth CE. Proposition 8 in nitanda2019gradient yields estimates for the Rademacher complexity and covering numbers.
Theorem 7 (Proposition 8 in nitanda2019gradient).
Suppose Assumptions (A1) and (A2) hold. Let $\forall w>0,\quad\forall m\in\mathbb{Z}_{+},\quad\forall T\in\mathbb{Z}_{+},\quad\forall\delta\in(0,1),\quad\text{and}\quad\forall S\text{ of size }n$. Then, there exists a universal constant $C>0$ such that with probability at least $1-\delta$ with respect to the initialization of $\Theta^{(0)}$, the empirical Rademacher complexity satisfies:
$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ $\leq Cm^{1/2-\beta}D_{w,m,T}(1+K_{1}+K_{2})$
$\qquad\cdot\sqrt{\frac{d}{n}\log\left(n(1+K_{1}+K_{2})\left(\log(m/\delta)+D_{w,m,T}^{2}\right)\right)}.$ (254)
where
$D_{w,m,T}$ $=\sqrt{wT}$ (255)
Moreover, when $\sigma$ is convex and $\sigma(0)=0$, with probability at least $1-\delta$ over a random initialization of $\theta^{(0)}$, we have:
$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ $\leq\frac{8K_{1}m^{1/2-\beta}}{\sqrt{n}}\left(D_{w,m,T}+\sqrt{\frac{\log(Am/\delta)}{b}}\right).$ (256)
By substituting Theorems 6 and 7 into the uniform convergence bound in Theorem 2, we obtain the following guarantee on the smooth CE for two-layer neural networks.
Corollary 7.
Suppose Assumption 1 and the settings of Theorem 6 hold. Then with probability $1-\delta$ over the random draw of $S_{\mathrm{tr}}$ and initial parameter $\theta_{0}$, we have
$\min_{t\in\{0,\dots,T-1\}}\mathrm{smCE}^{\sigma}(g_{\theta^{(t)}},\mu)\leq C_{1}\sqrt{\frac{1}{\gamma^{2}T}\left(\frac{m^{2\beta-1}}{w}+K\right)}+$ (257)
$\frac{C_{2}}{\sqrt{n}}+C_{3}m^{1/2-\beta}D_{w,m,T}(1+K_{1}+K_{2})\cdot$
$\qquad\cdot\sqrt{\frac{d}{n}\log\left(n(1+K_{1}+K_{2})\left(\log(2m/\delta)+D_{w,m,T}^{2}\right)\right)})+6\sqrt{\frac{\log\frac{6}{\delta}}{n}}.$ (258)
where $C_{1},\dots,C_{3}$ are universal constants.
Moreover, when $\sigma$ is convex and $\sigma(0)=0$, with probability at least $1-\delta$ over the random draw of $S_{\mathrm{tr}}$ and initial parameter $\theta_{0}$, such that we have:
$\min_{t\in\{0,\dots,T-1\}}\mathrm{smCE}^{\sigma}(g_{\theta^{(t)}},\mu)\leq C_{1}\sqrt{\frac{1}{\gamma^{2}T}\left(\frac{m^{2\beta-1}}{w}+K\right)}+$ (259)
$\frac{C_{2}}{\sqrt{n}}+\frac{C_{4}K_{1}m^{1/2-\beta}}{\sqrt{n}}\left(D_{w,m,T}+\sqrt{\frac{\log(2Am/\delta)}{b}}\right)+6\sqrt{\frac{\log\frac{6}{\delta}}{n}}.$ (260)
where $C_{4}$ is a universal constant.
Proof.
We use the following upper bound from Mohri: with probability $1-\delta/2$, we have
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})\leq\hat{\mathfrak{R}}_{S}(\mathcal{F})+\sqrt{\frac{\log\frac{2}{\delta}}{2n}}.$ (261)
Then from Theorem 2 with probability $1-\delta$, we have
$\sup_{f\in\mathcal{F}}|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ $\leq\frac{C_{2}}{\sqrt{n}}+4\hat{\mathfrak{R}}_{S}(\mathcal{F})+6\sqrt{\frac{\log\frac{3}{\delta}}{n}}.$ (262)
We then substitute Eq. equation 7 and equation 256 and taking the union bound, we obtain the result. ∎
For completeness, we include the result on the classification error, adapted from Theorem 4 in nitanda18a:
Theorem 8 (Theorem 4 in nitanda18a).
Suppose Assumption 1 and the settings of Theorem 6 hold. Fix $\forall\epsilon>0$. Then, with probability at least $1-3\delta$ over a random initialization and random choice of training dataset $S_{\mathrm{tr}}$, we have
$\min_{t\in\{0,\dots,T-1\}}\mathbb{P}_{(X,Y)\sim\nu}\left[Yg_{\theta^{(t)}}(X)\leq 0\right]\leq C_{5}(1+\exp(\epsilon))C_{w,m,T}+3\sqrt{\frac{\log(2/\delta)}{2n}}$
$\quad+C_{6}m^{1/2-\beta}D_{w,m,T}(1+K_{1}+K_{2})\cdot\sqrt{\frac{d}{n}\log\left(n(1+K_{1}+K_{2})\left(\log(m/\delta)+D_{w,m,T}^{2}\right)\right)},$ (263)
where
$C_{w,m,T}$ $=\gamma^{-1}T^{-1/2}\left(m^{\beta-1/2}w^{-1/2}+\sqrt{K}\right)$ (264)
and $C_{5}$ and $C_{6}$ are universal constants.
Moreover, when $\sigma$ is convex and $\sigma(0)=0$, we can avoid the dependence on the dimension $d$. With probability at least $1-3\delta$ over a random initialization and random choice of training dataset $S_{\mathrm{tr}}$,
$\min_{t\in\{0,\dots,T-1\}}\mathbb{P}_{(X,Y)\sim\nu}\left[Yg_{\theta^{(t)}}(X)\leq 0\right]$ $\leq C_{5}(1+\exp(\epsilon))C_{w,m,T}+3\sqrt{\frac{\log(2/\delta)}{2n}}$
$\quad+C_{7}K_{1}m^{1/2-\beta}\frac{1}{\epsilon\sqrt{n}}\left(D_{w,m,T}+\sqrt{\frac{\log(Am/\delta)}{b}}\right).$ (265)
where $C_{7}$ is a universal constant.
G.2 Formal statement for Corollary 5
We restate Corollary 5 with formal assumptions
Corollary 8.
Suppose Assumption 1 and the settings of Theorem 6 hold. If for any $\epsilon>0$, the hyperparameters satisfy one of the following:
(i) $\beta\in[0,1)$, $m=\Omega(\gamma^{\frac{-2}{1-\beta}}\epsilon^{\frac{-1}{1-\beta}})$, $T=\Omega(\gamma^{-2}\epsilon^{-2})$, $w=\Theta(\gamma^{-2}\epsilon^{-2}T^{-1}m^{2\beta-1})$, $n=\tilde{\Omega}(\gamma^{-2}\epsilon^{-4})$,
(ii) $\beta=0$, $m=\Theta\left(\gamma^{-2}\epsilon^{-3/2}\log(1/\epsilon)\right)$, $T=\Theta\left(\gamma^{-2}\epsilon^{-1}\log^{2}(1/\epsilon)\right)$, $w=\Theta(m^{-1})$, $n=\tilde{\Omega}(\epsilon^{-2})$, then with probability at least $1-\delta$, gradient descent with the stepsize $w$ finds a parameter $\theta^{(t)}$ satisfying $\mathrm{smCE}(g_{\theta^{(t)}},S_{\mathrm{tr}})\leq\epsilon$ and $\mathbb{P}_{(X,Y)\sim\nu}\left[Yg_{\theta^{(t)}}(X)\leq 0\right]\leq\epsilon$ within $T$ iterations.
Proof.
The guarantee for the misclassification rate is adapted from Corollary 3 in nitanda18a. As for the smooth CE, since the complexity terms in the misclassification bound (Theorem 8) and the smooth CE bound (Corollary 7) are of the same order, we also obtain a corresponding guarantee for the smooth CE. ∎
Appendix H Relationships of the proper scoring rule and functional gradient
In the main paper, we discussed the relationships involving the functional gradient only over the training dataset. Here, we provide the corresponding results for the population smooth CE.
We begin by connecting the post-processing gap with the functional gradient. First, we recall the definition of the functional gradient:
Definition 11.
Let $\mathcal{H}$ be a Hilbert space and $h$ be a function on $\mathcal{H}$. For $\xi\in\mathcal{H}$, we say that $h$ is Fréchet differentiable at $\xi$ in $\mathcal{H}$ if there exists an element $\nabla_{\xi}h(\xi)\in\mathcal{H}$ such that
$h(\zeta)=h(\xi)+\langle\nabla_{\xi}h(\xi),\zeta-\xi\rangle_{\mathcal{H}}+o(\|\xi-\zeta\|_{\mathcal{H}}).$ (266)
Moreover, for simplicity, we refer to $\nabla_{\xi}h(\xi)$ as the functional gradient.
See luenberger1997optimization for more details.
Next, we focus on the case of the squared loss. To simplify the notation, we express $r(f(x))=f(x)+h(f(x))$ where $h$ is the post-processing function in the definition of the smooth CE. Then we set $h\to\ell_{\mathrm{sq}}$, $\zeta\to f(x)$, and $\xi\to r(f(x))$ in the definition of the functional gradient. This results in the following relation.
$\ell_{\mathrm{sq}}(f(x),y)$ (267)
$=\ell_{\mathrm{sq}}(r(f(x)),y)-\nabla_{f}\ell_{\mathrm{sq}}(f(x),y)\cdot(r(f(x))-f(x))-\frac{1}{2}\|r(f(x))-f(x)\|^{2},$ (268)
where $\nabla_{f}\ell_{\mathrm{sq}}(f(x),y)$ denotes the partial derivative of $\ell_{\mathrm{sq}}(f,y)$ with respect to $f$. Thus, we observe that $\nabla_{f}\ell_{\mathrm{sq}}(f(x),y)$ corresponds to the functional gradient. Therefore,
$\mathrm{smCE}(f,\mathcal{D})^{2}\leq\mathrm{pGap}(f,\mathcal{D})$ $=\sup_{h}\mathbb{E}\left[-\nabla_{f}\ell_{\mathrm{sq}}(f(X),Y)\cdot h(f(X))-\frac{1}{2}\|h(f(X))\|^{2}\right]$ (269)
$\leq\sup_{h}\mathbb{E}\left[-\nabla_{f}\ell_{\mathrm{sq}}(f(X),Y)\cdot h(f(X))\right],$ (270)
where the supremum is taken over all 1-Lipschitz functions.
Next, for the cross-entropy loss, recall that $\nabla_{s}^{2}\ell^{\psi}(s,y)=p(1-p)$, where $p=\mathrm{pred}_{\psi}(s)=e^{s}/(1+e^{s})$. This implies that the cross-entropy loss is $1/4$-Lipschitz with respect to $s$. Therefore, by Taylor’s theorem, there exists $\tilde{p}\in(0,1/4]$ such that
$\ell^{\psi}(g(x),y)=\ell^{\psi}(r(g(x)),y)-\nabla_{g}\ell^{\psi}(g(x),y)\cdot(r(g(x))-g(x))-\frac{1}{2}\tilde{p}(1-\tilde{p})\|r(g(x))-g(x)\|^{2}.$ (271)
Thus, we obtain
$2\mathrm{smCE}^{(\sigma)}(g,\mathcal{D})^{2}$ $\leq\mathrm{pGap}^{(\psi,1/4)}(g,\mathcal{D})$ (272)
$=\sup_{h}\mathbb{E}\left[-\nabla_{g}\ell^{\psi}(g(X),Y)\cdot h(g(X))-\frac{1}{2}\tilde{p}(1-\tilde{p})\|h(g(X))\|^{2}\right]$ (273)
$\leq\sup_{h}\mathbb{E}\left[-\nabla_{g}\ell^{\psi}(g(X),Y)\cdot h(g(X))\right],$ (274)
where the supremum is taken over all $1/4$-Lipschitz functions.
From Eqs. equation 267 and equation 271, we obtain upper bounds on the training smooth CE in terms of the functional gradient evaluated on the training dataset:
$\mathrm{smCE}(f,S_{\mathrm{tr}})^{2}\leq\sup_{\eta}\frac{1}{n}\sum_{i=1}^{n}\left[-\nabla_{f}\ell_{\mathrm{sq}}(f(X_{i}),Y_{i})\cdot\eta(f(X_{i}))\right],$ (275)
and
$2\mathrm{smCE}^{(\psi,1/4)}(g,S_{\mathrm{tr}})^{2}\leq\sup_{\eta}\frac{1}{n}\sum_{i=1}^{n}\left[-\nabla_{g}\ell^{\psi}(g(X_{i}),Y_{i})\cdot\eta(g(X_{i}))\right].$ (276)
Finally, we remark on the case of general proper losses. As discussed in Appendix A, the post-processing gap can be defined more generally. If the loss is Fréchet differentiable, then similar relationships to those derived above for the cross-entropy loss can be obtained.
Appendix I Experiment
In this section, we numerically validate our theoretical findings regarding smooth CE for GBTs and two-layer neural networks presented in Section 4.
We evaluate the behavior of several metrics on both training and test datasets, including cross-entropy loss, accuracy, functional gradient norm (denoted as "Func Grad" in the experiments), and binning ECE, MMCE, and smooth CE. These evaluations are conducted by varying the number of iterations $T$ and the training dataset size $n$. Each experiment is repeated 10 times with different random seeds, and we report the mean and standard deviation.
For the binning ECE, we use 10 equally spaced bins. For MMCE, we use the Laplacian kernel with a bandwidth set to 1. Definitions and additional details of the calibration metrics are provided in Appendix B.
I.1 Gradient Boosting Trees
We conduct numerical experiments on the $L_{1}$ Gradient Boosting Tree with cross-entropy loss, as described in Algorithm 1. The main objective is to investigate how the training and test smooth CE behave on both toy and real datasets, as predicted by Theorem 3 and Corollary 2.
The toy dataset consists of two classes with equal probability, i.e.,
$P(Y=0)=P(Y=1)=\frac{1}{2}.$
Given the class label $Y\in\{0,1\}$, the conditional distribution of $X\in\mathbb{R}^{2}$ is:
[…]
These distributions have overlapping supports for $Y=1$ and $Y=0$.
First, we set $m=3$, which satisfies the condition $m\geq d$, and present the results in Figure 1. In Figure 1(a), we fix the training dataset size ($n=200$) and increase the number of boosting iterations $T$. The left panel shows the behavior of the metrics on the training dataset, the middle panel shows the same metrics on the test dataset, and the right panel reports the generalization gaps for both log loss and smooth CE. From the left panel, we observe that both the cross-entropy loss and smooth CE decrease as $T$ increases, consistent with Theorem 3. In contrast, the middle panel shows that neither metric necessarily decreases on the test dataset as $T$ increases. This is consistent with the well-known overfitting behavior of boosting, and demonstrates that smooth CE exhibits overfitting trends similar to those of the cross-entropy loss, as predicted by Corollary 2. The right panel confirms this overfitting phenomenon for both cross-entropy loss and smooth CE.
Next, we investigate how these metrics behave as the training sample size $n$ increases. Motivated by Corollary 2, we scale the number of boosting iterations $T$ proportionally to $\sqrt{n}$. The results are shown in Figure 1(b). We observe that the test calibration metrics decrease monotonically as $n$ increases, indicating improved generalization. These results show that GBTs, when trained with sufficient samples, can achieve both high accuracy and low smooth CE.
(a) Increasing the number of iterations $T$
(b) Increasing the number of training samples $n$
Figure 1: GBT experiments on the toy dataset defined in Eq. equation 278 with $m\geq d$
Next, we set $m=1$, which does not satisfy the condition $m\geq d$. The results are shown in Figure 2. Despite the violation of the assumption, we can see that the behavior of the calibration metrics seems similar to that of $m\geq d$. This indicates that there is a possibility that we might relax the assumptions of our theoretical analysis.
(a) Increasing the number of iterations $T$
(b) Increasing the number of training samples $n$
Figure 2: GBT experiments on the toy dataset defined in Eq. equation 278 with $m<d$
Next, using the UCI Breast Cancer dataset, we observed how the metrics behave on the training and test datasets by increasing the number of iterations $T$ while keeping the training dataset size fixed. The results are shown in Figure 3. Since $d=30$ in this dataset, we considered two settings: (i) $m=30$, which satisfies the assumption $m\geq d$, and (ii) $m=3$, which violates the assumption.
We found that the results closely resemble those of the toy dataset in Figure 1. Even when the assumption is violated, the training smooth CE decreases monotonically as the number of iterations increases. Combined with the results in Figure 2, this observation motivates us to consider relaxing the depth assumption in our theoretical analysis. We leave this for future work.
(a) Increasing the number of iterations $T$ ($m=30$, $m\geq d$)
(b) Increasing the number of iterations $T$ ($m=3$, $m<d$)
Figure 3: GBT experiments for UCI breast cancer dataset $d=30$
Finally, we also remark that the binning ECE, smooth CE, and MMCE exhibit similar behavior across all experimental settings. This observation is consistent with the discussion in Appendix B.
I.2 Numerical experiments over two-layer neural network
We conducted similar experiments using a two-layer neural network corresponding to Section 4.3. We prepared the toy dataset and designed the two-layer neural network so that Assumption 1 in Appendix G is satisfied. We used the sigmoid activation function, initialized the parameters independently from the standard Gaussian distribution, and set the number of hidden units to be even to satisfy Assumption (A3). The coefficients $a_{r}$ were also set according to the specification in Assumption (A3). Unless otherwise stated, we used 300 hidden units. For optimization, we used full-batch gradient descent (GD) with a fixed step size of $w=0.01$.
First, we performed experiments on the same toy dataset used for GBTs and evaluated the training and test metrics by increasing the number of GD iterations while fixing the training dataset size ($n=200$). Figure 4(a) reports the same set of metrics as in the GBT experiments (Figure 1). We observed that the smooth CE decreases monotonically, consistent with the upper bound behavior predicted by Eq. equation 27. However, from the middle and right panels of Figure 4(a), we observe that the test smooth CE does not decrease as the number of training iterations increases. We also confirmed that the generalization gap of the cross-entropy loss increases as $T$ increases. According to Eq. equation 257, the complexity term grows with the number of iterations, which may explain this phenomenon. On the other hand, we did not observe a clear increase in the generalization gap of the smooth CE as $T$ increased.
Next, we varied the training dataset size while setting the number of iterations to $T=\mathcal{O}(\sqrt{n})$, and evaluated various metrics. The results are shown in Figure 4(b). From the middle and right panels, we observe that both the test smooth CE and the generalization gap decrease monotonically as $n$ increases. This is consistent with the theoretical discussion provided in Appendix G.
(a) Increasing the number of iterations $T$
(b) Increasing the number of training samples $n$
Figure 4: Two-layer neural network using toydata defined in Eq. equation 278
The experiments in Figure 4 are based on a toy dataset where the separability condition is easily satisfied. Next, we consider a toy dataset in which the separability condition is more difficult to satisfy, constructed as follows. Assume that $n$ is even and define $n_{0}=n_{1}=\frac{n}{2}$. Let $\{Z_{i}\}_{i=1}^{n_{0}}\sim\mathcal{N}(0,\sigma^{2}I_{2})$ with $\sigma=0.05$ be shared random base vectors. Define two classes:
$X_{i}^{(0)}$ $=Z_{i}+\begin{bmatrix}0.1\\ 0.1\end{bmatrix}+\varepsilon^{(0)}_{i},\quad\varepsilon^{(0)}_{i}\sim\mathcal{N}(0,\tau^{2}I_{2}),$ (279)
$X_{i}^{(1)}$ $=-Z_{i}+\begin{bmatrix}-0.1\\ -0.1\end{bmatrix}+\varepsilon^{(1)}_{i},\quad\varepsilon^{(1)}_{i}\sim\mathcal{N}(0,\tau^{2}I_{2}),$ (280)
with $\tau=0.01$. Each $X_{i}^{(0)}$ is labeled $Y=0$, and each $X_{i}^{(1)}$ is labeled $Y=1$. Due to the symmetric structure of the data, this setting makes the margin assumption more difficult to satisfy than in the previous toy dataset. The results are shown in Figure 5, where we increase the training dataset size and evaluate the relevant statistics on both training and test datasets. Even under this more challenging setting, we observe that the test smooth CE decreases monotonically.
Figure 5: Two-layer neural network using toy dataset defined in Eq. equation 279
Next, we used the UCI Breast Cancer dataset and evaluated various statistics on both the training and test datasets, following the same procedure as in Figure 4(a). Here, we varied the number of hidden units from a relatively small size (10) to a larger value ($100$). The results are shown in Figure 6.
Consistent with the results in Figure 4(a), we observed that the training statistics decrease monotonically in both small and large hidden unit settings. However, the test statistics did not show any corresponding improvement.
(a) Increasing the number of iterations $T$ when the hidden unit size is 10
(b) Increasing the number of iterations $T$ when the hidden unit size is 100
Figure 6: UCI breast cancer dataset $d=30$